### v2

# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

Abstract
**footnotetext: Equal Contribution22footnotetext: Work done during internship at Adobe Research
Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language processing. However, due to sequential processing through a stack of transformer layers, autoregressive decoding faces significant computation/latency challenges, particularly in resource-constrained environments like mobile and edge devices. Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors - 1) Early exit, and 2) Input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence. Both the above strategies have limitations - the former cannot be applied to handle KV Caching necessary for speed-ups in modern framework and the latter does not capture the variation in layer importance across tasks or more generally, across input sequences. To address both limitations, we propose FiRST, an algorithm that reduces inference latency by using layer-specific routers to select a subset of transformer layers adaptively for each input sequence - the prompt (during the prefill stage) decides which layers will be skipped during decoding. FiRST preserves compatibility with KV caching enabling faster inference while being quality-aware. FiRST is model-agnostic and can be easily enabled on any pre-trained LLM. Our approach reveals that input adaptivity is critical - indeed, different task-specific middle layers play a crucial role in evolving hidden representations depending on tasks. Extensive experiments show that FiRST significantly reduces latency while outperforming other layer selection strategies in quality metics. It retains competitive performance to base model (without layer skipping) and in some cases, even improves upon it. FiRST is thus a promising and efficient solution for LLM deployment in low-resource environments.
1 Introduction
Large Language Models (LLM’s) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks. However, the scale of these LLM’s characterized by billions of parameters hinder their adoption in resource-constrained environments with memory, latency and compute being the main challenges. In this work, we focus on the latency aspect which becomes the most significant bottleneck for tasks such as machine translation, question answering, summarization particularly on devices, such as laptops and mobile phones. As noted by (Schuster et al., 2022), the auto-regressive nature of decoding in LLM’s further pronounces the latency bottleneck.
Transformer based LLMs have several stacks of layers (including attention and FFN layers) leading to high latency and compute requirements, making inference very slow or even infeasible in resource constrained settings. This is because of the sequential processing of tokens through all the layers for every input sequence and task. However, it is important to note that in the real world, there is a lot of heterogeneity in input sequences and tasks. (Schuster et al., 2022; Sun et al., 2022) noted that the generations made by LLMs can have varying levels of difficulty and certain generations can be solved with reduced compute, by exiting the transformer stack early. At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation. Motivated by these and other related works, we hypothesize that different sequential combinations of layers are important for different input sequences and tasks. Learning the right sequential combination of layers can help reduce inference latency and compute for on-device scenarios. However, there are several challenges. Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling and be learnable with minimal compute/training overhead.
In the last few years, several promising approaches have been proposed in literature that adaptively prune layers at each decoding step. Token-level early exit proposed in (Schuster et al., 2022; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level. (Elhoushi et al., 2024; Elbayad et al., 2020; Zhang et al., 2019) extended this idea to incorporate layer skipping at a token level during training. While token level early exit is a useful idea in theory, it suffers from a major limitation of incompatible KV caching in practice (Del Corro et al., 2023). The incompatibility stems from having to recompute KV caches for preceding tokens if we have a delayed exit point for latter tokens, often resulting in loss of early exit advantages. This limits its practical adoption since KV cache is crucial in significantly speeding up auto-regressive decoding.
Recently, (Liu et al., 2024; Del Corro et al., 2023; Song et al., 2024) have proposed input-agnostic layer skipping at token level, that handle KV cache appropriately as well as retain the advantage of adaptive partial computation. In these solutions, tokens exit at pre-determined layers irrespective of the input sequence, and for all sequences in a batch, tokens at the same position in a sequence exit at the same layer. Furthermore, tokens at latter parts of the sequence are constrained to exit earlier than the previous tokens to ensure that there is no redundant KV cache re-computation. These solutions are heuristic based and impose hard rules and constraints irrespective of input sequences, which can lead to drop in output quality. Others (Jaiswal et al., 2024; Chen et al., 2024) have proposed skipping layers by identifying redundant ones through computing cosine similarity of (input/output) representations of a layer. However, their strategy does not take into account that several middle layers are crucial (see (Liu et al., 2024)) and furthermore, final prediction capability of full model is not taken into account while deciding which layers to skip. Importantly, in none of the works described above, the strategy of selecting layers for skipping is sequence dependent. Furthermore, they do not consider finetuning the models in a way such that not only the performance improves but also the model learns to skip layers appropriately.
Our goal is to design an input-adaptive learnable layer selection strategy with quality aware latency gains that is also able to handle the KV cache appropriately. Ideally, for every input sequence and task, we want to predict the optimal (sequential) combination of layers at inference time, such that quality loss is minimum and the latency gains are as high as possible. We want to do this with expending very little compute/additional training and appropriate handling of KV cache. We propose to do this via training routers. Based on the output of each layer for a sequence, a router will decide whether or not to skip the subsequent layer in the transformer architecture. Since the decision is at a sequence level, KV cache issues do not arise, as all tokens in a sequence would pass through the same set of layers. Finally, we fine-tune the model combined with trained routers using LoRA adapters to improve the quality significantly while retaining the latency gains. As an added bonus, LoRA finetuning smoothens the layer skipping and further highlights the varied importance of layers based on input sequence. We summarize our contributions:
  1. 1.
    We propose a training and inference algorithm FiRST that incorporates layer-specific routers for selecting layers in an input-adaptive manner. The layer selection is uniform for all tokens in a sequence, thus handling KV caches without introducing additional compute/latency. FiRST can be applied on top of any pre-trained model.
  2. 2.
    We propose LoRA based finetuning on top of router based layer selection to improve quality while retaining latency gains. This also smoothens layer selection.
  3. 3.
    Finally, we demonstrate an extensive set of experiments with FiRST on multiple datasets for 3 distinct tasks namely Machine Translation, Summarization and Question Answering, and 2 different open model architectures namely LLaMA-3-8B and LLaMA-3.2-3B. We show that for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines and prior works.
Due to space constraints, we delegate a study of other related works and orthogonal approaches (for e.g. model compression) for exploring latency/performance tradeoff to Appendix A.1.
2 Problem Statement
Our goal is to exploit the heterogeneity in inputs and tasks to selectively use LLM layers in a quality-aware manner for reducing inference latency and compute for on-device constraints. Ideally, we want to select an optimal sub-sequence of layers within a transformer architecture for a given input and task, such that the overall latency, as well as expended computation, are both low, while quality is comparable to the un-modified case where every input sequence passes through every layer. For ease of explanation, without loss of generality, we assume the task is same and simply consider an input sequence for describing the problem.
Let us consider an an input sequence $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ with $n$ tokens. Let there be $m$ transformer layers in the model, where the $i^{th}$ transformer layer is represented as the function $\phi_{i}()$. As stated lucidly in (Wendler et al., 2024), $\mathcal{X}$ is first converted to an initial latent representation $\mathcal{H}_{0}=\{H^{1}_{0},H^{2}_{0},\ldots,H^{n}_{0}\}$, where $H^{0}_{j}\in\mathbb{R}^{D},\forall j\in[n]$ is a look-up from a learned embedding dictionary corresponding to the $j^{th}$ token. Thereafter, every transformer layer $\phi_{i}()$ operates on the latent vectors $\mathcal{H}_{i}$ to generate the embedding for the $i^{th}$ layer as follows. For the $j^{th}$ token,
$H^{j}_{i}$ $=H^{j}_{i-1}+\phi_{i}(H^{1}_{i-1},H^{2}_{i-1},\ldots,H^{j}_{i-1})$ (1)
Let the (gold) output or generated sequence for an input sequence $\mathcal{X}$ that passed through all $m$ layers of the model with full computation be $\mathcal{Y}^{*}_{\mathcal{X}}$. Our hypothesis is that for a given input sequence (and task), there exists an optimal subsequence of functions $\mathcal{F}_{OPT}(\mathcal{X})$ out of the full sequence $\{\phi_{i},i\in[m]\}$ such that the output generated by passing through this subsequence: $\mathcal{Y}_{OPT,\mathcal{X}}\approx\mathcal{Y}^{*}_{\mathcal{X}}$. More formally, if $Q$ is a quantitative quality measure on $\mathcal{Y}$, and $\epsilon\rightarrow 0$ is tolerance in deviation in quality from the gold output, then we hypothesize that there exists an optimal subsequence, using the minimum number of layers, $\mathcal{F}_{OPT}(\mathcal{X})$, such that:
$Q\left(\mathcal{Y}_{OPT,\mathcal{X}}\right)$ $\geq(1-\epsilon)Q\left(\mathcal{Y}^{*}_{\mathcal{X}}\right),\forall\mathcal{X}.$ (2)
The optimality above is with respect to the minimum subsequence of layers that can help achieve the above, to minimize latency while keeping quality unaffected. Note that, the optimal subsequence $\mathcal{F}_{OPT}(\mathcal{X})$ need to be obey the same autoregressive computation on previous tokens as given in Equation 1. Hence, any algorithm that determines the optimal subsequence, need to be compatible with KV cache handling, to avoid the re-computation of values for tokens preceding the current token.
The potential number of subsequences for $m$ layers is $2^{m}$, hence a brute force is infeasible and also beats the purpose of such a layer selection in the first place: reducing latency and compute. In the absence of any known substructure in the behaviour of the latent layers on each input sequence, it is difficult to arrive at the optimal solution polynomially or with low additional latency or compute.
We propose to learn an approximation of the optimal subsequence of layers for any input sequence with low additional latency and minimal training.
3 Proposed Solution: FiRST
Let us first understand what it entails to learn an optimal subsequence of layers for any input. Consider the full transformer sequence to be =$\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$. Any optimal subsequence for an input $\mathcal{X}$: $\mathcal{F}_{OPT,\mathcal{X}}$ could be thought of as finding an optimal path through a binary tree of functions. Formally, let every level in the binary tree correspond to a transformer layer and the $0^{th}$ layer corresponds to the initial embedding look up; i.e., at depth $i\in[m]$, there would be $2^{i}$ nodes, each corresponding to either $\phi_{i}$ or $\overline{\phi_{i}}$, where the former denotes that a particular transformer layer is included in the optimal path whereas the latter denotes that it is not included. Each (of the $2^{i-1}$ nodes) $\phi_{i}$ or $\overline{\phi_{i}}$ has two children, corresponding to the next transformer layer: $\phi_{i+1}$ and $\overline{\phi_{i+1}}$ (See Figure 1). In such a tree structure, for example, the path $\{\phi_{i},\overline{\phi_{i+1}},\phi_{i+2}\}$ indicates the subsequence of transformer layers $\{\phi_{i},\phi_{i+2}\}$. For any transformer layer $\phi_{i}$ in this tree, let $Anc(\phi_{i})=k,0\leq k<i$ denote the the lowest ancestor node where the corresponding transformer node $\phi_{k}$ is included in the sequence. In the above example, $Anc(\phi_{i+2})=\phi_{i}$.
Figure 1: Binary Tree representation of layer selection.
Consider a sequence of functions $\mathcal{F}$, where for level $i$, $Anc(\phi_{i})=\phi_{k}$. The autoregressive computations for the $j^{th}$ token in the input sequence (originally Eq 1), would now be modified as:
$H^{j}_{i}$ =$=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
Our problem translates to navigating this binary tree to find the optimal path $\mathcal{F}_{OPT}$ for an input sequence and task. Since there are $2^{m}$ paths in this tree, we propose to approximate the optimal by making a decision in a greedy fashion at each node. Formally, we add a (lightweight and fast) router $R_{i}$ before every transformer layer $\phi_{i}$ in the model, that will predict whether $\phi_{i}$ will be selected or not.
Our aim is to learn to predict the layer choice at a sequence level (not token) to maintain compatibility with the autoregressive computations and avoid re-computation of of KV cache values. Moreover, we should spend minimal compute for learning the $R_{i}$ functions. Finally, $R_{i}$ functions should not add any significant latency to the overall computation.
FiRST modifies any off-shelf pre-trained transformer based model by incorporating and training a router or probability function $R_{i}$ before every transformer layer $\phi_{i}$. The output of $R_{i}$ is a score $\rho_{i}$ denoting the probability of selecting $\phi_{i}$ in the layer sequence. During inference, $\rho_{i}$ is rounded to determine selection of $\phi_{i}$. Let $\lfloor\rho_{i}\rceil=1\text{ if }\rho_{i}\geq 0.5,\text{ else }0$. Equation 1 is now modified as:
$H^{j}_{i}$ $=H^{j}_{i-1}+\lfloor\rho_{i}\rceil\cdot\phi_{i}(H^{1}_{i-1},H^{2}_{i-1},\ldots,H^{j}_{i-1})$
This recursively approximates Eq 3 for the optimal $\mathcal{F}$ in a probabilistic, greedy manner. We train the functions $R_{i}$ on datasets and tasks, and further fine tune using LoRA adapters to make the layer selections smooth and improve the output quality.
4 FiRST Framework and Algorithm
In this section, we describe the training and inference frameworks for FiRST in details. We discuss how to train Routers to be adaptive to input sequences. Given an off-the-shelf pre-trained LLM, we propose two training phases. In the first phase, we train a router for each layer that decides whether the input sequence should skip the layer. In the second phase, to tackle the issue of unseen skipping during pre-training, we fine-tune the router-augmented LLM keeping router weights fixed to ensure the model improves performance on the target dataset without reducing the skipping level.
4.1 Adaptive Router Module
The adaptive router module is a single-layer neural network without bias, positioned before every layer in the model. During training of the router, all model parameters except the router weights remain frozen. For the first layer, it takes the tokenized input, and for each of the subsequent layers, it takes the output of the preceding layer as input. Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer $(i-1)$ and outputs a $B\times n\times 1$ tensor. Subsequently, corresponding to each value (or, token) in the $B\times n\times 1$ tensor, we apply a sigmoid function to ensure that all entries in the tensor are in the interval $[0,1]$. Following this, we take a mean operation at the sequence level - we take a mean of all the weights in a sequence to output a $B\times 1\times 1$ tensor. For each sequence in the batch, the corresponding entry is the probability $\rho_{i}$ with which the sequence passes through the layer $i$. The input sequence skips the layer $i$ with probability $1-\rho_{i}$. During training, the output of a layer is modified using a skip connection, incorporating the probability $\rho_{i}$ (see Figure: 2).
The routers are trained to encourage skipping by reducing the probabilities $\{\rho_{i}\}_{i}$ using a regularizer, to approximate the optimal subsequence for minimizing the latency. The training task is modeled as a language modeling task, specifically next token prediction. The loss function comprises of 3 terms:
  • Cross-entropy loss: Standard difference between actual and predicted probability distributions to ensure the quality of generation: $\mathcal{L}_{\mathrm{CE}}=-\sum_{x\in\mathcal{X}}\mathcal{Y}^{*}_{\mathcal{X}}\log(\hat{\mathcal{Y}})$.
  • Regularization loss: Adds a penalty term to reduce overfitting to noise: $\mathcal{L}_{\mathrm{Reg}}=\sum_{i\in[m]}||R_{i}||^{2}$, where $||R_{i}||^{2}$ denotes the $\ell_{2}$ norm of the router weights for the $i^{\text{th}}$ layer router, and there are $m$ layers in the model.
  • Non-skip penalization loss: This is the summation of probability values across all layers of the model architecture: $\mathcal{L}_{\mathrm{PP}}=\sum_{i\in[m]}\rho_{i}$
The total loss $\mathcal{L}$ is a linear combination of these three terms: $\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\lambda\cdot\mathcal{L}_{\mathrm{Reg}}+\alpha\cdot\mathcal{L}_{\mathrm{PP}}$, where $\alpha$ manages the tradeoff between quality and latency.
4.2 LoRA Compensation Module
Skipping layers naturally leads to some performance loss - especially so since the pre-trained model was not trained to skip layers. To compensate for the loss in performance caused by skipping layers, we finetune the router-augmented pre-trained model on the downstream task 111similar to Quantization Aware Training such as QLoRA Dettmers et al. (2024) - compensates for model compression using Low Rank Adapters (LoRA). During finetuning, the router parameters are frozen while trainable LoRA adapters are added to both the FFN (Feed-Forward Network) and the attention modules of each layer of the pre-trained model. In order to maintain the skipping level, we again add a non-skip penalization loss component during finetuning with scaling hyper-parameter $\beta$. This is essential even though the router weights are frozen because standard finetuning alters the hidden representations of the input sequence in a manner such that no layers are skipped. Note that the LoRA adapters do not lead to any latency overhead during inference.
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93 56.94 35.56 82.66
Base 37.17 18.57 87.13 38.02 22.46 68.95
15 Skip Decode Router + LoRA 23.04 10.52 55.62 28.73 15.84 55.98
Router 3.99 1.18 23.33 4.74 2.33 21.75
Random Skip Router + LoRA 30.43 10.98 66.25 47.88 25.75 67.32
Router 26.54 8.77 60.27 36.60 18.66 59.89
Unified Skip Router + LoRA 28.92 10.64 59.34 46.61 25.01 69.58
Router 23.23 7.85 59.26 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 38.01 17.89 82.14 48.35 26.57 68.63
Router 28.83 11.8 67.74 17.55 8.68 42.76
25 Skip Decode Router + LoRA 13.67 6.00 31.47 20.03 10.85 33.85
Router 3.24 0.92 21.55 3.78 1.84 20.93
Random Skip Router + LoRA 6.01 0.91 29.71 11.69 4.66 27.73
Router 3.65 0.49 29.95 7.37 2.81 35.16
Unified Skip Router + LoRA 15.67 3.36 31.69 34.90 15.75 50.59
Router 12.58 2.65 32.15 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 17.84 4.14 34.95 35.79 15.66 56.92
Router 9.67 1.37 26.01 11.01 3.23 25.45
Table 1: Machine Translation Results for LLaMA-3-8B: BLEU-1, BLEU-2, and COMET scores are reported for English-to-German and English-to-Chinese tasks across varying skip levels. FiRST consistently achieves the highest BLEU-1 and BLEU-2 at 15% skipping, for both translation directions. Similarly, high BLEU-1 and COMET scores are obtained for 25% skipping rate.
4.3 Inference for FiRST
During inference, for the input sequence, each router (corresponding to a layer) outputs a number in the interval $[0,1]$. If this number is greater than or equal to 0.5, the sequence passes through the layer. Otherwise, the sequence skips the layer (Fig. 2).
Figure 2: Skip connection used for router training. With probability $p$, the sequence is processed by the layer and with probability $1-p$, the layer is skipped. During inference, routers make the decision of whether a sequence will skip a particular layer or pass through it.
Below, we discuss some salient points about the functioning of the router during inference to handle KV Cache appropriately:
  1. 1.
    Prefill phase handling: Skipping is not allowed during prefill phase. This ensures the first token is generated correctly, which is crucial for WMT tasks, as they are highly sensitive to the correct generation of the first token in the target language. It has been observed in prior works (Liu et al., 2024) that skipping during prefill phase is detrimental to performance during inference.
  2. 2.
    Fixed router decisions during decoding and handling KV Cache: During the prefill phase, the decisions made by the routers are cached. During the decoding phase, every token adheres to the cached decision made during prefill. In other words, for a particular layer, if a router outputs a number less than 0.5 during prefill, the number is fixed for the decoding steps and therefore the same layer will be skipped by all tokens during decoding. Similarly, if the router outputs a number more than 0.5 during prefill, the same layer will be processing all tokens during decoding. Such a step ensures that for each decoding step and each layer that is not skipped, the KV cache for all previous tokens is available for that layer. This approach effectively addresses the caching issues encountered in early exit strategies, ensuring consistent decisions across the decoding process.
5 Experiments
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72 51.81 30.04 79.13
Base 31.19 13.67 81.66 32.10 17.92 61.84
15 Skip Decode Router + LoRA 23.24 11.07 44.58 38.14 21.68 46.70
Router 9.93 3.89 32.74 9.59 4.84 34.14
Random Skip Router + LoRA 18.99 4.78 47.26 38.75 17.39 57.79
Router 12.11 2.77 36.30 13.56 5.64 35.86
Unified Skip Router + LoRA 21.65 5.93 44.72 36.96 17.35 57.10
Router 16.41 3.99 39.81 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 24.68 7.13 60.29 45.66 23.66 67.45
Router 16.47 4.11 43.04 22.69 10.92 54.55
25 Skip Decode Router + LoRA 16.85 7.62 32.33 27.81 15.74 42.01
Router 9.14 3.51 27.64 7.01 3.41 29.51
Random Skip Router + LoRA 8.95 1.07 30.97 25.12 9.85 45.53
Router 5.32 0.69 27.22 10.07 3.69 31.50
Unified Skip Router + LoRA 15.14 2.65 39.81 30.56 12.27 42.30
Router 10.21 1.62 30.86 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 18.89 4.06 45.38 32.92 13.74 41.66
Router 9.51 1.44 29.08 10.46 3.55 27.83
Table 2: Machine Translation Results for LLaMA-3.2-3B: BLEU-1, BLEU-2, and COMET scores are shown for English-to-German and English-to-Chinese tasks at different skip levels. FiRST consistently achieves the highest BLEU-1 and COMET scores for English-to-German at both skipping percentages and highest BLEU-1 for English-to-Chinese.
Skip (%) Model Type EM F1
0 Original Model Base + LoRA 73.93 85.99
Base 19.46 36.73
10 Skip Decode R + LoRA 60.14 65.33
Router 16.38 31.48
Random Skip R + LoRA 65.73 80.08
Router 18.25 33.75
Unified Skip R + LoRA 55.54 74.58
Router 17.39 32.91
FiRST (Ours) R + LoRA 70.85 83.61
Router 14.58 31.52
20 Skip Decode R + LoRA 45.00 55.10
Router 10.68 26.69
Random Skip R + LoRA 47.79 66.37
Router 6.71 22.46
Unified Skip R + LoRA 52.87 69.28
Router 18.18 32.51
FiRST (Ours) R + LoRA 60.60 75.49
Router 13.21 27.48
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.07 84.17
Base 18.92 37.74
10 Skip Decode R + LoRA 60.79 75.00
Router 20.00 31.55
Random Skip R + LoRA 64.78 77.27
Router 13.76 28.59
Unified Skip R + LoRA 65.03 77.53
Router 13.16 32.31
FiRST R + LoRA 69.44 81.35
Router 12.79 28.37
20 Skip Decode R + LoRA 40.12 40.00
Router 20.45 37.62
Random Skip R + LoRA 11.32 38.34
Router 6.75 15.51
Unified Skip R + LoRA 37.39 52.49
Router 7.81 18.20
FiRST R + LoRA 39.70 54.59
Router 5.52 15.33
Table 3: Quality Analysis on Question Answering (SQuAD v1.1): LLaMA-3-8B (left) and LLaMA-3.2-3B (right). Exact Match (EM) and F1 scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.87 28.46 16.99
Base 82.29 23.49 14.66
15 Skip Decode R + LoRA 84.74 22.04 17.54
Router 82.53 13.68 9.30
Random Skip R + LoRA 83.70 24.60 15.01
Router 81.10 19.64 13.07
Unified Skip R + LoRA 84.25 24.35 14.3
Router 80.3 16.61 10.95
FiRST (Ours) R + LoRA 85.14 31.8 20.13
Router 81.25 20.2 13.01
27 Skip Decode R + LoRA 79.92 10.67 10.32
Router 77.27 9.59 7.00
Random Skip R + LoRA 76.40 11.45 7.89
Router 77.45 12.56 9.08
Unified Skip R + LoRA 80.28 15.94 9.89
Router 77.43 10.97 7.68
FiRST (Ours) R + LoRA 77.5 14.65 10.45
Router 75.6 9.39 6.92
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.89 28.37 17.02
Base 71.85 19.34 12.00
15 Skip Decode R + LoRA 83.20 21.71 13.74
Router 80.97 9.74 6.87
Random Skip R + LoRA 79.52 20.18 12.10
Router 68.20 10.10 7.10
Unified Skip R + LoRA 81.53 18.89 11.72
Router 70.01 12.49 8.68
FiRST (Ours) R + LoRA 83.17 26.47 16.79
Router 70.98 16.47 10.51
24 Skip Decode R + LoRA 78.55 15.83 6.74
Router 76.91 13.29 8.86
Random Skip R + LoRA 80.00 16.33 10.07
Router 67.88 8.49 5.96
Unified Skip R + LoRA 79.31 15.88 10.69
Router 68.86 9.17 6.97
FiRST (Ours) R + LoRA 80.25 21.28 13.89
Router 69.17 12.36 8.43
Table 4: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
We conduct experiments on three benchmark tasks namely Machine Translation, Text Summarization and Question Answering demonstrating both robustness and scalability of FiRST.
Datasets: For machine translation, we use WMT development sets (2017–2020) for English-to-Chinese and English-to-German tasks, evaluating performance on the WMT 2022 test set, which covers diverse domains such as news, social media, e-commerce, and conversational contexts. For summarization, we use the CNN/DailyMail dataset, with 4000 randomly selected training samples and evaluation on the standard test set of 11,490 samples. In Question Answering, we utilize SQuAD v1.0, training on 4000 randomly selected samples from the 100k question-answer pairs and evaluating on the validation set, as test set labels are unavailable. Appendix A.2 contains detailed descriptions.
Evaluation Metrics: We use standard metrics to evaluate the quality of generated output in each of the three tasks. For Machine Translation, we benchmark using BLEU scores and COMET where the latter provides a more nuanced assessment beyond $n$-gram used in BLEU. For Summarization, we use ROUGE scores and BERT Score - as before, the latter captures meaning-based similarity. For Question Answering, we use Exact Match and F1 Score as the metrics to benchmark the output quality. Finally, for benchmarking latency, we look at the TPOT (Time Per Output Token). TPOT evaluates the average time taken to produce each output token and is calculated for GPU to gauge overall decoding performance. Appendix A.3 contains a detailed description of evaluation metrics. Additionally, hyper-parameters used during training and inference can be found in Appendix A.4.
5.1 Baselines for comparison
We report the latency improvement and quality numbers relative to the base models (no skipping).
  • Random Skipping: We skip a set of $k$ layers randomly where $k$ depends on the target speedup.
  • Skip Decode: We implement Skip Decode Del Corro et al. (2023) method that features a monotonic decrease in processing layers, enabling later tokens to leverage the computational resources used for earlier ones.
  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method that relies on using a heuristic-based strategy for retaining layers at fixed intervals. We replicate the algorithm in (Liu et al., 2024) and compare performance both with and without LoRA fine-tuning across various skipping percentages.
5.2 Experimental Results on Different Tasks
WMT: Our experiments demonstrate that FiRST performs consistently well in different machine translation tasks (EN-to-DE, EN-to-ZH) across different models. For LLaMA-3.8B, for $15\%$ skipping, FiRST achieves a latency improvement of 10-12% on TPOT (see Tables 1 and 5) while being competitive to the base model + LoRA (gold reference) in quality. In most cases, it significantly outperforms other layer skipping strategies (Random, Skip Decode, Unified Skipping) and in other cases, it is comparable in quality. In comparison to the gold output (Base + LoRA), FiRST is $\geq 80\%$ in COMET scores (semantic metric), $\geq 85\%$ in BLEU-1 scores, and $\geq 75\%$ in BLEU-2 scores (syntax metrics) for both EN-to-DE and EN-to-ZH translations. For $25\%$ skipping, FiRST achieves significant improvement in quality over other strategies, in almost all metrics, while achieving $\sim 18\%$ reduction in TPOT.
For LLaMA-3.2-3B, the latency improvement is $\sim 10\%$, with quality scores being significantly higher than other layer skipping strategies (Table 6). It is within $65-85\%$ of BLEU-1 and COMET scores (Table 2) for EN-DE and EN-ZH of the gold (LoRA fine-tuned base model).
CNN/DailyMail Dataset: For LLaMA-3-8B, at roughly 15% skipping level, our method outperforms the base model + LoRA (Table: 4) while obtaining a 12% improvement in TPOT (Table 5). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.88x
R + LoRA 25 0.82x 0.83x
R + LoRA 35 0.69x 0.68x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 20 0.81x
R + LoRA 27 0.76x
Table 5: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.91x
R + LoRA 25 0.78x 0.75x
R + LoRA 35 0.69x 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 24 0.79x
R + LoRA 28 0.77x
Table 6: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.83x
R + LoRA 28 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.84x
R + LoRA 28 0.72x
Table 7: TPOT variation of LLaMA-3-8B (left) and LLaMA-3.2-3B (right) on SQuAD dataset for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
SQuAD Dataset: For the LLaMA-3-8B model, FiRST is $>95\%$ in output quality of gold (base + LoRA) (Table 3), with overall latency gains of 6-16% (Table 7). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 3) with gains in latency of 6-16% overall (Table 7) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics. Detailed results for an additional skipping percentage are provided in Appendix A.8.
Layer-wise Skipping Patterns: Layer-wise skipping varies significantly across tasks, reflecting the task-specific importance of each layer. For LLaMA-3-8B at a 15% skipping rate, layers 7–9 and 21 are fully skipped in English-to-German, with partial skipping in layer 18. In English-to-Chinese, layers 7–9, 16, and 21 are fully skipped, while layer 20 is partially skipped. For Summarization, layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19, 21, and 26. Some layers are skipped less than 10%, indicating their necessity for specific sequences. This task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped, and skipping patterns depend on the input. Detailed statistics are in Appendix A.7.
Computational overhead of Routers: We also examine the computational overhead introduced by training the routers. This overhead remains minimal, as the router parameters are significantly smaller than the total number of trainable parameters in each model (0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B). The proportion of the total computational time spent on the router operations compared to the entire forward pass of the model is close to 0.3% for both models suggesting their overhead does not significantly affect the efficiency of the model during inference or training.
6 Conclusion
We propose a new framework FiRST for layer selection corresponding to input sequence and task towards reducing latency in a quality aware manner. This is sequence dependent and operates in a KV cache compatible manner. For an optimal skipping rate of around 15%, FiRST achieves 10-20% reduction in latency while being quality neutral (approximately 80% or more in quality metrics compared to base model) on multiple tasks/model architectures such as Machine Translation, Summarization and Question Answering, on well known open source datasets, and some cases, even improving upon base model, while significantly outperfoming other layer selection strategies on most quality metrics.
7 Limitations
FiRST algorithm selects layers in a greedy, myopic way one layer at a time, corresponding to sequences (and tasks). A more optimal way of doing this would be to estimate a subsequence of layers to traverse through instead of one layer at a time. We intend to address this in future work. We would like to select a more optimal subset (subsequence) of layers which will increase the output quality while reducing latency even further.
8 Ethical Concerns
There are no ethical concerns to the best of our knowledge.
Appendix A Appendix
A.1 Related Work
Early Exit:
Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022) where adaptive compute is used for different parts of the token sequence. While these approaches have been popular for encoder-only models which processes the entire sequence as a whole, they have faced challenges in generation tasks. The main limitation of these set of techniques are their inability to handle KV caching appropriately which is crucial for multi-fold speed-ups in current LLM architectures. We emphasize that in our work, we assign varying compute to sequences in different batches but within the same sequence, we assign the same compute to every token.
Input Agnostic Heuristics:
In Skip Decoding (Del Corro et al., 2023), initial tokens pass through more layers than later ones, contradicting the observation that later tokens are harder to decode (Liu et al., 2024). Additionally, Skip Decoding skips several bottom layers for most tokens, causing undesirable sub-network imbalance. To address this, Unified Layer Skipping (Liu et al., 2024) proposes a discrete skipping strategy that is uniform for all tokens in a sequence. Based on a latency budget, retained layer ids are passed through by all tokens, ensuring KV Cache handling and retaining key layers. However, the limitation of this approach is that skipping is independent of the input sequence. In contrast, early exit strategies adapt layer skipping to the input sequence, offering more flexibility. In (Fan et al., 2019), a method akin to dropout randomly skips layers during training, but this leads to performance decline during the pre-fill stage. FFN-SkipLLM (Jaiswal et al., 2024) constrains skipping to FFN layers to avoid KV Cache issues but fails to fully exploit redundancy as discussed already. Song et al. (2024) is a very recent work that also explores greedily identifying layers to skip while preserving the model performance on a calibration dataset - however there are two major limitations of this work which are resolved in our paper - (A) first of all, the layer selection strategy is sequence independent although it can be made task-dependent by calibrating on task-specific datapoints - our approach for skipping layers is sequence dependent and is based on the input to a layer (B) SLEB does not explore the impact of fine-tuning on the layers to be skipped. On the other hand, our skipping strategy incorporates the trained router already - intuitively, the knowledge of skipping is transferred during finetuning. Chen et al. (2024) is another recent work that compresses models by identifying redundant layers - this is done by computing the average similarity between input/output pairs of a layer. However, as outlined in Song et al. (2024), such an approach suffers from the limitation that it does not take into account the joint association between the layers while skipping multiple layers. Moreover, like (Jaiswal et al., 2024), this work is neither sequence dependent nor takes the final model predictions into account while identifying the layers to skip.
Model Compression and Quantization Aware Training:
Orthogonal approaches to explore the latency/memory-performance trade-off in Large Language Models aim to build smaller models that approximate the performance of larger ones with reduced memory and latency costs. Key techniques include: 1) compressing model parameters into fewer bits (Frantar et al., 2022; Lin et al., 2024; Lee et al., 2024; Saha et al., 2023); 2) pruning the network by removing components like attention heads or neurons based on heuristics (Frantar & Alistarh, 2023; Ma et al., 2023b); and 3) distilling the large model into a smaller, faster counterpart (Agarwal et al., 2023; Gu et al., 2024). For further details, we refer to the survey by (Zhu et al., 2023). A significant body of work (Dettmers et al., 2024; Liu et al., 2023b; Peri et al., 2020; Li et al., 2023) has focused on quantization-aware training to reduce memory footprints and mitigate performance loss, starting with QLoRA (Dettmers et al., 2024). In a similar vein, our work proposes fine-tuning router-augmented models to improve layer skipping and reduce performance degradation, as pre-trained models do not account for layer skipping, leading to higher degradation with vanilla skipping.
Network Pruning:
Another orthogonal approach to improve the inference speed-up is to prune redundant network weights by zeroing them out. There has been a significant body of work on pruning model weights Frantar et al. (2022); Frantar & Alistarh (2023); Sun et al. (2022); Zhang et al. (2023) - most of these works can be categorized into two clusters namely unstructured pruning and structured pruning. In case of unstructured pruning, there is no structure to the inserted zeros and achieving speedups with modern GPU hardware tailored towards dense matrix multiplication is challenging. In fact, more than 90% sparsity is typically required to achieve any significant speedup Wang (2020); Shi et al. (2020). Therefore, structured pruning which is more amenable to GPU hardware has become prominent (2:4 pruning and sub-channel pruning). However, realizing desired speedups through these techniques have been difficult Song et al. (2024). Moreover, several approaches for dynamically deleting entire rows or columns of weight matrices have been proposed Ma et al. (2023a); Ashkboos et al. (2024); Liu et al. (2023c) to retain dense matrices but two limitations remain - (A) hardware support is extremely limited for realizing speedup gains (B) extensive finetuning is necessary to align the sparsification with linguistic abilities - this is because, such pruning techniques were not observed by the model during pre-training. Finally, note that several prior works Tang et al. (2024); Oren et al. (2024); Xiao et al. (2023) have imposed (query aware/ query agnostic) sparsity in the KV cache matrices to speed up self-attention mechanism via clever selection of the critical tokens necessary from the KV cache.
A.2 Details of Datasets
Machine Translation: For translation tasks, namely English-to-Chinese and English-to-German, we employ the WMT development sets from 2017 to 2020 for training/fine-tuning following the methodology outlined in previous studies (Liu et al., 2023a; Jiao et al., 2023). Translation performance is evaluated using the test set from the WMT 2022 dataset (Kocmi et al., 2022) which was developed using recent content from diverse domains. These domains include news, social media, e-commerce, and conversational contexts.
(Details in Appendix: A.5, Table: 8). Summarization: We use the popular CNN-DailyMail (CNN/DM) (Hermann et al., 2015) dataset which is a large collection (over 300k) of text summarization pairs, created from CNN and Daily Mail news articles. Each datapoint in this dataset comprises of an article (the body of the news article with 683 words on average) and the corresponding highlights (article summary as written by the article author). While the training set contains more than 287k samples, we have randomly chosen 4k samples for training both routers and LoRA. During training in our framework, the number of trainable parameters is small in both phases - therefore a small subset of data points is sufficient for training.
Inference is performed on the standard test set with 11,490 samples.
Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.0) (Rajpurkar et al., 2016), a widely-used benchmark for machine Question Answering. The dataset consists of over 100k question-answer pairs posed by crowd-workers on a set of over 500 Wikipedia articles. Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage) and the corresponding answer (a text span from the corresponding reading passage). Similarly to the CNN/DM dataset, 4k samples are chosen as random to train both routers and LoRA. The training and validation split contains 87,599 and 10,570 samples respectively. Evaluation is performed on the validation set Schuster et al. (2022) as the test set labels are not publicly released.
A.3 Evaluation Metrics
Quality-Based Metrics for Translation task:
  • BLEU Score: BLEU (Bilingual Evaluation Understudy) scores are used to measure the quality of translations. BLEU compares n-grams of the candidate translation to n-grams of the reference translation, providing a score between 0 and 1, with higher scores indicating better translations. In this evaluation, NLTK BLEU is employed, focusing on BLEU-1 and BLEU-2 scores.
  • COMET: COMET (Cross-lingual Optimized Metric for Evaluation of Translation) is used to assess translation quality further. COMET evaluates translations using a model trained to correlate well with human judgments. Specifically, Unbabel/XCOMET-XL  [^2] https://github.com/Unbabel/COMET is used in this evaluation. COMET provides a more nuanced assessment of translation quality by considering the intricacies of both source and target languages, beyond the n-gram matching used in BLEU.
Quality based Metrics for Summarization Task:
  • BERTScore: This metric quantifies semantic similarity between texts by leveraging contextual word embeddings.
    BERTScore captures meaning-based similarity rather than relying on exact word matches, providing a nuanced evaluation of text generation quality.
  • ROUGE: (Recall-Oriented Understudy for Gisting Evaluation) is a common metric - ROUGE-1 refers to overlap of unigrams between the system summary and reference summary. Similarly, ROUGE-L measures longest matching sequence of words.
Quality based Metrics for Question Answering Task:
  • Exact Match: This metric measures the percentage of predictions that exactly match the ground truth answer.
  • F1 score: Since EM is a highly stringent metric, we also report the F1 score which provides a more flexible evaluation of answer prediction. This metric also takes into account near-matches.
A.4 Training and Inference Setup
  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively. Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization. The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps. The learning rate is set between $1e^{-4}$ and $3e^{-4}$ a cosine scheduler is used to adjust the learning rate. Gradients are accumulated after 5 steps and the regularization coefficient $\lambda$ is fixed at 0.01. For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor (lora_alpha) of 32. We set the non-skipping penalization regularizer hyper-parameter $\beta=\alpha/3$ after tuning to maintain skipping level - where $\alpha$ is the non-skipping penalization hyperparameter used in the first phase and set according to the target speedup ratio. For translation, the maximum sequence length is set to 128 for router training and 256 for LoRA training. Similarly, for summarization, the maximum sequence length is set to 500 and 700 respectively. For Question Answering, this length is set to 512. Prompts for the different tasks regarding training/inference are shown in Appendix A.6.
  • Inference settings: For all the tasks, we set the temperature to 0.8 and enable top-k sampling over 10 tokens. The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD. Caching is turned on during inference.
A.5 Training and testing split
WMT Summarization RC
Eng-to-German Eng-to-Chinese CNN/DM SQuAD
Train 3505 8983 3400 3400
Validation 876 998 600 600
Test 2038 2038 11490 10570
Table 8: Train-Validation-Test split for WMT, CNN/DM and SQuAD datasets
Figure 3: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on English-to-German Machine Translation Task. The graph shows how different layers contribute to the skipping behavior for the same dataset. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 4: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on English-to-Chinese Machine Translation Task. The graph shows how different layers contribute to the skipping behavior for the same dataset. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 5: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on CNN Summarization Task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 6: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on SQuAD Question-Answering Task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
A.6 Prompt Details
The prompt structures used for both training and inference are as follows:
  • For the machine translation task (English-to-German or English-to-Chinese), the following general prompt structure is used to train the routers and during final inference:
    ### Instruction:
        Translate the following sentences from English
    to German.
    ### Input:
        {Text to be translated}
    ### Response:
  • For the summarization task (used in CNN/DailyMail dataset), the prompt structure used is:
    ### Instruction:
        Summarize the news article in around 100-200
    words.
    ### Input:
        {Article to be summarized}
    ### Response:
  • For the Question Answering task (used in SQuAD dataset), the following prompt structure is utilized:
    ### Instruction:
        Answer the question based on the given passage.
    ### Passage:
        {context}
    ### Question:
        {Question to be answered}
    ### Response:
[…]
A.7 Layer-wise Skipping Statistics
Table 15 present the fraction of sequences that skip a particular block during the task for the LLaMA-3-8B model. If the corresponding cell in a row shows a value of 80.00, it implies that 80% of the sequences skip this block. It is important to note that the decision regarding which block to skip varies across different datasets and tasks. Additionally, partial skipping in some blocks, with varying percentages, suggests that while some sequences consider the layer important, others do not and therefore skip it during the decoding phase.
Figures 3, 4, 5, and 6 depict the distribution of skipped blocks when the model is configured to skip approximately 15% of the layers. These plots demonstrate the variation in block skipping based on the task and dataset, highlighting the differences in block importance. The bar plots further show that the determination of important layers depends not only on the dataset but also on the model architecture. For instance, Figure 3 illustrates that layers 7 and 9 are entirely skipped in LLaMA-3-8b, while they are rarely skipped in LLaMA-3.2-3b.
A.8 Detailed Result Table
Tables 9 and 11 present the detailed results for the LLaMA-3.2-3B model, while Tables 10 and 12 summarize the performance of the LLaMA-3-8B model for an additional skipping percentage on Machine Translation Task. The results are reported using BLEU (BLEU-1, BLEU-2) and COMET metrics, highlighting performance across different skipping percentages. Similarly, Table 13 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L. Lastly, Table 14 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72
Base 31.19 13.67 81.66
15 Skip Decode Router + LoRA 23.24 11.07 44.58
Router 9.93 3.89 32.74
Random Skip Router + LoRA 18.99 4.78 47.26
Router 12.11 2.77 36.30
Unified Skip Router + LoRA 21.65 5.93 44.72
Router 16.41 3.99 39.81
FiRST (Ours) Router + LoRA 24.68 7.13 60.29
Router 16.47 4.11 43.04
25 Skip Decode Router + LoRA 16.85 7.62 32.33
Router 9.14 3.51 27.64
Random Skip Router + LoRA 8.95 1.07 30.97
Router 5.32 0.69 27.22
Unified Skip Router + LoRA 15.14 2.65 39.81
Router 10.21 1.62 30.86
FiRST (Ours) Router + LoRA 18.89 4.06 45.38
Router 9.51 1.44 29.08
35 Skip Decode Router + LoRA 1.62 0.36 23.30
Router 1.71 0.27 19.34
Random Skip Router + LoRA 4.72 0.15 25.08
Router 1.96 0.06 21.18
Unified Skip Router + LoRA 1.18 0.05 19.65
Router 0.90 0.03 20.36
FiRST (Ours) Router + LoRA 9.47 1.29 27.45
Router 5.74 0.56 25.03
Table 9: Machine Translation Results for English to German on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93
Base 37.17 18.57 87.13
15 Skip Decode Router + LoRA 23.04 10.52 55.62
Router 3.99 1.18 23.33
Random Skip Router + LoRA 30.43 10.98 66.25
Router 26.54 8.77 60.27
Unified Skip Router + LoRA 28.92 10.64 59.34
Router 23.23 7.85 59.26
FiRST (Ours) Router + LoRA 38.01 17.89 82.14
Router 28.83 11.8 67.74
25 Skip Decode Router + LoRA 13.67 6.00 31.47
Router 3.24 0.92 21.55
Random Skip Router + LoRA 6.01 0.91 29.71
Router 3.65 0.49 29.95
Unified Skip Router + LoRA 15.67 3.36 31.69
Router 12.58 2.65 32.15
FiRST (Ours) Router + LoRA 17.84 4.14 34.95
Router 9.67 1.37 26.01
35 Skip Decode Router + LoRA 5.55 0.33 23.85
Router 3.03 0.82 20.03
Random Skip Router + LoRA 1.80 0.12 25.56
Router 1.44 0.06 25.34
Unified Skip Router + LoRA 6.44 0.77 22.05
Router 3.92 0.51 22.88
FiRST (Ours) Router + LoRA 6.39 0.42 19.96
Router 3.7 0.14 21.41
Table 10: Machine Translation Results for English to German on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 51.81 30.04 79.13
Base 32.10 17.92 61.84
15 Skip Decode Router + LoRA 38.14 21.68 46.70
Router 9.59 4.84 34.14
Random Skip Router + LoRA 38.75 17.39 57.79
Router 13.56 5.64 35.86
Unified Skip Router + LoRA 36.96 17.35 57.10
Router 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 45.66 23.66 67.45
Router 22.69 10.92 54.55
25 Skip Decode Router + LoRA 27.81 15.74 42.01
Router 7.01 3.41 29.51
Random Skip Router + LoRA 25.12 9.85 45.53
Router 10.07 3.69 31.50
Unified Skip Router + LoRA 30.56 12.27 42.30
Router 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 32.92 13.74 41.66
Router 10.46 3.55 27.83
35 Skip Decode Router + LoRA 2.84 1.10 22.30
Router 2.36 1.04 18.32
Random Skip Router + LoRA 4.17 1.52 32.81
Router 2.11 0.76 23.63
Unified Skip Router + LoRA 3.26 0.52 19.12
Router 2.20 0.37 19.82
FiRST (Ours) Router + LoRA 20.01 6.96 28.10
Router 6.54 1.78 23.13
Table 11: Machine Translation Results for English to Chinese on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 56.94 35.56 82.66
Base 38.02 22.46 68.95
15 Skip Decode Router + LoRA 28.73 15.84 55.98
Router 4.74 2.33 21.75
Random Skip Router + LoRA 47.88 25.75 67.32
Router 36.60 18.66 59.89
Unified Skip Router + LoRA 46.61 25.01 69.58
Router 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 48.35 26.57 68.63
Router 17.55 8.68 42.76
25 Skip Decode Router + LoRA 20.03 10.85 33.85
Router 3.78 1.84 20.93
Random Skip Router + LoRA 11.69 4.66 27.73
Router 7.37 2.81 35.16
Unified Skip Router + LoRA 34.90 15.75 50.59
Router 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 35.79 15.66 56.92
Router 11.01 3.23 25.45
35 Skip Decode Router + LoRA 12.24 6.27 25.23
Router 3.64 1.74 22.84
Random Skip Router + LoRA 7.38 2.19 27.35
Router 4.47 1.12 29.46
Unified Skip Router + LoRA 7.51 2.10 20.25
Router 3.87 1.06 21.24
FiRST (Ours) Router + LoRA 15.66 3.95 26.80
Router 6.13 1.54 22.89
Table 12: Machine Translation Results for English to Chinese on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.87 28.46 16.99
Base 82.29 23.49 14.66
15 Skip Decode R + LoRA 84.74 22.04 17.54
Router 82.53 13.68 9.30
Random Skip R + LoRA 83.70 24.60 15.01
Router 81.10 19.64 13.07
Unified Skip R + LoRA 84.25 24.35 14.3
Router 80.3 16.61 10.95
FiRST (Ours) R + LoRA 85.14 31.8 20.13
Router 81.25 20.2 13.01
20 Skip Decode R + LoRA 82.57 20.41 14.87
Router 81.62 13.48 9.19
Random Skip R + LoRA 81.39 21.57 13.83
Router 79.23 15.51 10.93
Unified Skip R + LoRA 82.93 22.3 13.37
Router 80.32 16.51 11.15
FiRST (Ours) R + LoRA 82.8 27.65 17.84
Router 79.32 16.28 10.85
27 Skip Decode R + LoRA 79.92 10.67 10.32
Router 77.27 9.59 7.00
Random Skip R + LoRA 76.40 11.45 7.89
Router 77.45 12.56 9.08
Unified Skip R + LoRA 80.28 15.94 9.89
Router 77.43 10.97 7.68
FiRST (Ours) R + LoRA 77.5 14.65 10.45
Router 75.6 9.39 6.92
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.89 28.37 17.02
Base 71.85 19.34 12.00
15 Skip Decode R + LoRA 83.20 21.71 13.74
Router 80.97 9.74 6.87
Random Skip R + LoRA 79.52 20.18 12.10
Router 68.20 10.10 7.10
Unified Skip R + LoRA 81.53 18.89 11.72
Router 70.01 12.49 8.68
FiRST (Ours) R + LoRA 83.17 26.47 16.79
Router 70.98 16.47 10.51
24 Skip Decode R + LoRA 78.55 15.83 6.74
Router 76.91 13.29 8.86
Random Skip R + LoRA 80.00 16.33 10.07
Router 67.88 8.49 5.96
Unified Skip R + LoRA 79.31 15.88 10.69
Router 68.86 9.17 6.97
FiRST (Ours) R + LoRA 80.25 21.28 13.89
Router 69.17 12.36 8.43
28 Skip Decode R + LoRA 70.69 8.76 6.74
Router 40.99 2.05 1.23
Random Skip R + LoRA 79.45 14.69 9.23
Router 67.48 8.14 5.64
Unified Skip R + LoRA 78.57 11.74 7.47
Router 68.23 8.12 5.66
FiRST (Ours) R + LoRA 77.48 15.98 11.14
Router 67.14 8.09 6.00
Table 13: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning. FiRST with fine-tuning, improves upon Unified Skipping for all skipping levels on both Rouge-1 and Rouge-L and is competitive on BERT F1.
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.93 85.99
Base 19.46 36.73
10 Skip Decode R + LoRA 60.14 65.33
Router 16.38 31.48
Random Skip R + LoRA 65.73 80.08
Router 18.25 33.75
Unified Skip R + LoRA 55.54 74.58
Router 17.39 32.91
FiRST (Ours) R + LoRA 70.85 83.61
Router 14.58 31.52
20 Skip Decode R + LoRA 45.00 55.10
Router 10.68 26.69
Random Skip R + LoRA 47.79 66.37
Router 6.71 22.46
Unified Skip R + LoRA 52.87 69.28
Router 18.18 32.51
FiRST (Ours) R + LoRA 60.60 75.49
Router 13.21 27.48
30 Skip Decode R + LoRA 30.77 48.38
Router 10.67 28.52
Random Skip R + LoRA 25.45 42.68
Router 3.55 15.49
Unified Skip R + LoRA 25.61 38.55
Router 15.19 28.11
FiRST (Ours) R + LoRA 38.20 52.68
Router 3.64 13.30
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.07 84.17
Base 18.92 37.74
10 Skip Decode R + LoRA 60.79 75.00
Router 20.00 31.55
Random Skip R + LoRA 64.78 77.27
Router 13.76 28.59
Unified Skip R + LoRA 65.03 77.53
Router 13.16 32.31
FiRST R + LoRA 69.44 81.35
Router 12.79 28.37
20 Skip Decode R + LoRA 40.12 40.00
Router 20.45 37.62
Random Skip R + LoRA 11.32 38.34
Router 6.75 15.51
Unified Skip R + LoRA 37.39 52.49
Router 7.81 18.20
FiRST R + LoRA 39.70 54.59
Router 5.52 15.33
30 Skip Decode R + LoRA 20.68 23.08
Router 0.78 3.83
Random Skip R + LoRA 0.30 8.87
Router 0.62 5.88
Unified Skip R + LoRA 7.39 13.72
Router 0.37 6.15
FiRST R + LoRA 33.99 50.37
Router 2.55 10.26
Table 14: SQuAD performance on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): EM (Exact Match) and F1 scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
Layer $\downarrow$ R R+L R R+L R R+L
$\alpha\rightarrow$ 0.005 0.005 0.01 0.01 0.025 0.025
0 0.00 0.00 0.00 0.00 0.00 0.00
1 0.00 0.00 0.00 0.00 0.00 0.00
2 0.00 0.00 0.00 0.00 0.00 0.00
3 0.00 0.00 0.00 0.00 0.00 0.00
4 0.00 0.00 0.00 0.00 0.00 0.00
5 0.00 0.00 0.00 0.00 0.00 0.00
6 0.00 0.00 0.00 0.00 0.00 0.00
7 100.00 100.00 100.00 100.00 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00
9 100.00 100.00 0.00 0.00 0.10 1.32
10 0.00 0.00 0.00 0.00 89.25 72.67
11 0.00 0.00 0.00 0.00 0.00 0.00
12 0.00 0.00 100.00 100.00 100.00 100.00
13 0.00 0.00 0.00 0.00 0.00 0.00
14 0.00 0.00 0.00 0.59 0.00 0.25
15 0.00 0.00 100.00 99.95 100.00 99.36
16 0.00 0.00 0.10 2.45 100.00 99.95
17 0.00 0.00 0.00 0.00 0.00 0.00
18 97.79 32.24 100.00 100.00 100.00 100.00
19 0.00 0.00 99.85 91.17 100.00 99.61
20 0.00 0.00 100.00 100.00 100.00 99.46
21 99.85 98.72 100.00 100.00 100.00 100.00
22 0.00 0.00 0.00 0.00 0.00 0.00
23 0.00 0.00 24.14 0.79 99.75 91.66
24 0.00 0.00 0.00 0.00 0.00 0.00
25 0.00 0.00 0.00 0.00 0.00 0.00
26 0.00 0.00 57.31 2.45 100.00 86.02
27 0.00 0.00 0.00 0.00 0.00 0.00
28 0.00 0.00 0.00 0.00 0.00 0.05
29 0.00 0.00 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00 0.00 0.00
31 0.00 0.00 0.00 0.00 0.00 0.00
Avg 15.55 13.47 27.54 24.92 37.16 35.95
Table 15: Variation in skipping percentage (15-35%) with the Non-skip Penalization Loss coefficient $\alpha$ for LLaMA-3-8B on Machine Translation (English-to-German). As $\alpha$ increases, the skipping percentage also increases for both the Router-Only and Router-Augmented LoRA fine-tuned models. Similar trends are observed for other datasets as well.

### v3

@@
 **footnotetext: Equal Contribution22footnotetext: Work done during internship at Adobe Research
-Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language processing.
-However, due to sequential processing through a stack of transformer layers, autoregressive decoding faces significant computation/latency challenges, particularly in resource-constrained environments like mobile and edge devices.
-Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors - 1) Early exit, and 2) Input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence.
-Both the above strategies have limitations - the former cannot be applied to handle KV Caching necessary for speed-ups in modern framework and the latter does not capture the variation in layer importance across tasks or more generally, across input sequences.
-To address both limitations, we propose FiRST, an algorithm that reduces inference latency by using layer-specific routers to select a subset of transformer layers adaptively for each input sequence - the prompt (during the prefill stage) decides which layers will be skipped during decoding.
-FiRST preserves compatibility with KV caching enabling faster inference while being quality-aware.
-FiRST is model-agnostic and can be easily enabled on any pre-trained LLM.
-Our approach reveals that input adaptivity is critical - indeed, different task-specific middle layers play a crucial role in evolving hidden representations depending on tasks.
-Extensive experiments show that FiRST significantly reduces latency while outperforming other layer selection strategies in quality metics.
-It retains competitive performance to base model (without layer skipping) and in some cases, even improves upon it.
+Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language tasks.
+However, due to sequential processing through multiple transformer layers, autoregressive decoding faces significant computational challenges, particularly in resource-constrained environments like mobile and edge devices.
+Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors: (1) early exit, and (2) input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence.
+Both the above strategies have limitations—the former cannot be applied in the presence of KV caching, which is essential for speed-ups in modern inference frameworks, and the latter fails to capture variation in layer importance across tasks or, more generally, across input sequences.
+To address these limitations, we propose FiRST, a model-agnostic framework that reduces inference latency by using layer-specific routers to adaptively skip transformer layers during decoding, based on routing decisions made from the input prompt in the prefill stage.
+FiRST remains fully compatible with KV caching, enabling faster decoding while maintaining quality.
+Our method reveals that input adaptivity is essential—different tasks rely on different subsets of layers to evolve meaningful representations.
+Extensive experiments show that FiRST significantly reduces latency while outperforming existing layer selection strategies in quality.
+It retains performance comparable to the base model without skipping.
 FiRST is thus a promising and efficient solution for LLM deployment in low-resource environments.
 1 Introduction
-Large Language Models (LLM’s) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks.
-However, the scale of these LLM’s characterized by billions of parameters hinder their adoption in resource-constrained environments with memory, latency and compute being the main challenges.
-In this work, we focus on the latency aspect which becomes the most significant bottleneck for tasks such as machine translation, question answering, summarization particularly on devices, such as laptops and mobile phones.
-As noted by (Schuster et al., 2022), the auto-regressive nature of decoding in LLM’s further pronounces the latency bottleneck.
-Transformer based LLMs have several stacks of layers (including attention and FFN layers) leading to high latency and compute requirements, making inference very slow or even infeasible in resource constrained settings.
-This is because of the sequential processing of tokens through all the layers for every input sequence and task.
+Large Language Models (LLMs) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks.
+However, the massive scale of LLMs—often involving billions of parameters—poses significant challenges for deployment in resource-constrained environments, where memory, compute, and especially latency become critical bottlenecks.
+In this work, we focus on addressing the latency issue, which is particularly pronounced in edge settings such as laptops and mobile devices.
+As noted by Schuster et al.
+(2022a), the auto-regressive nature of decoding in LLMs further amplifies this bottleneck.
+Transformer-based LLMs consist of multiple stacked layers—including attention and feed-forward networks—which result in high latency and computational cost.
+This makes inference slow or even impractical in resource-constrained environments.
+The inefficiency stems from the need to process each input token sequentially through all layers, irrespective of the input sequence or task.
 However, it is important to note that in the real world, there is a lot of heterogeneity in input sequences and tasks.
-(Schuster et al., 2022; Sun et al., 2022) noted that the generations made by LLMs can have varying levels of difficulty and certain generations can be solved with reduced compute, by exiting the transformer stack early.
-At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation.
+(Schuster et al., 2022a; Sun et al., 2022) observed that LLM generations can have varying levels of difficulty and certain generations can be solved with reduced computation, by exiting the transformer stack early.
+At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer-based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation.
 Motivated by these and other related works, we hypothesize that different sequential combinations of layers are important for different input sequences and tasks.
@@
 However, there are several challenges.
-Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling and be learnable with minimal compute/training overhead.
+Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling, and be learnable with minimal compute/training overhead.
 In the last few years, several promising approaches have been proposed in literature that adaptively prune layers at each decoding step.
-Token-level early exit proposed in (Schuster et al., 2022; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level.
+Token-level early exit proposed in (Schuster et al., 2022a; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level.
 (Elhoushi et al., 2024; Elbayad et al., 2020; Zhang et al., 2019) extended this idea to incorporate layer skipping at a token level during training.
@@
 Importantly, in none of the works described above, the strategy of selecting layers for skipping is sequence dependent.
-Furthermore, they do not consider finetuning the models in a way such that not only the performance improves but also the model learns to skip layers appropriately.
-Our goal is to design an input-adaptive learnable layer selection strategy with quality aware latency gains that is also able to handle the KV cache appropriately.
-Ideally, for every input sequence and task, we want to predict the optimal (sequential) combination of layers at inference time, such that quality loss is minimum and the latency gains are as high as possible.
-We want to do this with expending very little compute/additional training and appropriate handling of KV cache.
-We propose to do this via training routers.
-Based on the output of each layer for a sequence, a router will decide whether or not to skip the subsequent layer in the transformer architecture.
-Since the decision is at a sequence level, KV cache issues do not arise, as all tokens in a sequence would pass through the same set of layers.
-Finally, we fine-tune the model combined with trained routers using LoRA adapters to improve the quality significantly while retaining the latency gains.
-As an added bonus, LoRA finetuning smoothens the layer skipping and further highlights the varied importance of layers based on input sequence.
-We summarize our contributions:
+Furthermore, they do not consider fine-tuning the models in a way such that not only the performance improves but the model also learns to skip layers appropriately.
+Due to space constraints, we delegate a study of other related works and orthogonal approaches (for e.g. model compression) for exploring latency/performance tradeoff to Appendix A.1.
+Our goal is to design an input-adaptive, learnable layer selection strategy that provides quality-aware latency improvements while properly handling KV caching.
+Ideally, for each input sequence and task, we want to predict a sequential combination of layers to run during inference—minimizing quality loss while achieving as much latency gain as possible.
+We also want to do this with minimal computational overhead or extra training.
+To achieve this, we propose training routers.
+At each layer, the router looks at the current sequence representation and decides whether to skip the next layer.
+Since the decision is made at the sequence level, all tokens follow the same path through the model, avoiding KV cache inconsistencies during decoding.Finally, we fine-tune the model with trained routers using LoRA adapters to recover any quality drop introduced by layer skipping, while preserving latency gains.
+LoRA fine-tuning also smoothens layer skipping and further highlights the varied importance of layers based on input sequence.
+Our key contributions include:
   1. 1.
-    We propose a training and inference algorithm FiRST that incorporates layer-specific routers for selecting layers in an input-adaptive manner.
-The layer selection is uniform for all tokens in a sequence, thus handling KV caches without introducing additional compute/latency.
-FiRST can be applied on top of any pre-trained model.
+    We propose FiRST, a training and inference algorithm that uses layer-specific routers to perform input-adaptive layer selection.
+All tokens in a sequence follow the same selected layers, ensuring compatibility with KV caching and avoiding additional compute or latency overhead.
+FiRST is model-agnostic and can be applied on top of any pre-trained LLM.
   2. 2.
-    We propose LoRA based finetuning on top of router based layer selection to improve quality while retaining latency gains. This also smoothens layer selection.
+    We introduce a LoRA-based fine-tuning approach on top of router-based layer selection to recover quality while maintaining latency gains.
+This also encourages smoother and a more stable layer selection.
   3. 3.
-    Finally, we demonstrate an extensive set of experiments with FiRST on multiple datasets for 3 distinct tasks namely Machine Translation, Summarization and Question Answering, and 2 different open model architectures namely LLaMA-3-8B and LLaMA-3.2-3B.
-We show that for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines and prior works.
-Due to space constraints, we delegate a study of other related works and orthogonal approaches (for e.g. model compression) for exploring latency/performance tradeoff to Appendix A.1.
+    Finally, we conduct extensive experiments with FiRST across multiple datasets spanning three distinct tasks: Machine Translation, Summarization, and Question Answering.
+We evaluate on two open-source model architectures—LLaMA-3-8B and LLaMA-3.2-3B—and show that, for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines.
 2 Problem Statement
@@
 Thereafter, every transformer layer $\phi_{i}()$ operates on the latent vectors $\mathcal{H}_{i}$ to generate the embedding for the $i^{th}$ layer as follows.
-For the $j^{th}$ token,
+For the $j^{th}$ token, the embedding at layer $i$ is computed as follows:
 $H^{j}_{i}$ $=H^{j}_{i-1}+\phi_{i}(H^{1}_{i-1},H^{2}_{i-1},\ldots,H^{j}_{i-1})$ (1)
@@
 In the second phase, to tackle the issue of unseen skipping during pre-training, we fine-tune the router-augmented LLM keeping router weights fixed to ensure the model improves performance on the target dataset without reducing the skipping level.
+While joint training of the router and LoRA modules is theoretically possible, we find it introduces optimization instability (see Appendix A.4 for further discussion on this design choice and its impact).
 4.1 Adaptive Router Module
@@
 Note that the LoRA adapters do not lead to any latency overhead during inference.
-Skip (%) Model Type English-to-German English-to-Chinese
-BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 41.78 21.74 93 56.94 35.56 82.66
-Base 37.17 18.57 87.13 38.02 22.46 68.95
-15 Skip Decode Router + LoRA 23.04 10.52 55.62 28.73 15.84 55.98
-Router 3.99 1.18 23.33 4.74 2.33 21.75
-Random Skip Router + LoRA 30.43 10.98 66.25 47.88 25.75 67.32
-Router 26.54 8.77 60.27 36.60 18.66 59.89
-Unified Skip Router + LoRA 28.92 10.64 59.34 46.61 25.01 69.58
-Router 23.23 7.85 59.26 27.28 13.35 54.57
-FiRST (Ours) Router + LoRA 38.01 17.89 82.14 48.35 26.57 68.63
-Router 28.83 11.8 67.74 17.55 8.68 42.76
-25 Skip Decode Router + LoRA 13.67 6.00 31.47 20.03 10.85 33.85
-Router 3.24 0.92 21.55 3.78 1.84 20.93
-Random Skip Router + LoRA 6.01 0.91 29.71 11.69 4.66 27.73
-Router 3.65 0.49 29.95 7.37 2.81 35.16
-Unified Skip Router + LoRA 15.67 3.36 31.69 34.90 15.75 50.59
-Router 12.58 2.65 32.15 17.74 7.35 38.74
-FiRST (Ours) Router + LoRA 17.84 4.14 34.95 35.79 15.66 56.92
-Router 9.67 1.37 26.01 11.01 3.23 25.45
-Table 1: Machine Translation Results for LLaMA-3-8B: BLEU-1, BLEU-2, and COMET scores are reported for English-to-German and English-to-Chinese tasks across varying skip levels.
-FiRST consistently achieves the highest BLEU-1 and BLEU-2 at 15% skipping, for both translation directions.
-Similarly, high BLEU-1 and COMET scores are obtained for 25% skipping rate.
+En-to-De En-to-Zh
+Skip (%) Model Type LLaMA-3-8B LLaMA-3.2-3B LLaMA-3-8B LLaMA-3.2-3B
+BLEU COMET BLEU COMET BLEU COMET BLEU COMET
+0 Original Model Base + LoRA 0.199 93.00 0.160 89.72 0.333 82.66 0.278 79.13
+Base 0.169 87.13 0.125 81.66 0.208 68.95 0.166 61.84
+15 Skip Decode Router + LoRA 0.094 55.62 0.105 44.58 0.149 55.98 0.2 46.7
+Router 0.019 23.33 0.055 32.74 0.03 21.75 0.062 34.14
+Random Skip Router + LoRA 0.097 66.25 0.079 47.26 0.237 67.32 0.154 57.79
+Router 0.132 60.27 0.048 36.3 0.168 59.89 0.077 35.86
+Unified Skip Router + LoRA 0.153 59.34 0.095 44.72 0.23 69.58 0.157 57.1
+Router 0.117 59.26 0.067 39.81 0.122 54.57 0.087 45.16
+FiRST (Ours) Router + LoRA 0.161 82.14 0.113 60.29 0.247 68.63 0.218 67.45
+Router 0.108 67.74 0.069 43.04 0.08 42.76 0.1 54.55
+25 Skip Decode Router + LoRA 0.07 31.47 0.077 32.33 0.088 33.85 0.147 42.01
+Router 0.015 21.55 0.05 27.64 0.024 20.93 0.045 29.51
+Random Skip Router + LoRA 0.018 29.71 0.023 30.97 0.065 27.73 0.137 45.53
+Router 0.011 29.95 0.015 27.22 0.04 35.16 0.053 31.5
+Unified Skip Router + LoRA 0.06 31.69 0.05 39.81 0.142 50.59 0.108 42.3
+Router 0.048 32.15 0.032 30.86 0.068 38.74 0.079 31.63
+FiRST (Ours) Router + LoRA 0.071 34.95 0.072 45.38 0.119 56.92 0.126 41.66
+Router 0.029 26.01 0.029 29.08 0.051 25.45 0.053 27.83
+Table 1: Machine Translation results for the English-to-German and English-to-Chinese task for both models: BLEU and COMET scores are reported across varying skip levels.
+FiRST (Ours) performs consistently well across both translation directions.
 4.3 Inference for FiRST
@@
 5 Experiments
-Skip (%) Model Type English-to-German English-to-Chinese
-BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 37.57 17.43 89.72 51.81 30.04 79.13
-Base 31.19 13.67 81.66 32.10 17.92 61.84
-15 Skip Decode Router + LoRA 23.24 11.07 44.58 38.14 21.68 46.70
-Router 9.93 3.89 32.74 9.59 4.84 34.14
-Random Skip Router + LoRA 18.99 4.78 47.26 38.75 17.39 57.79
-Router 12.11 2.77 36.30 13.56 5.64 35.86
-Unified Skip Router + LoRA 21.65 5.93 44.72 36.96 17.35 57.10
-Router 16.41 3.99 39.81 22.43 9.37 45.16
-FiRST (Ours) Router + LoRA 24.68 7.13 60.29 45.66 23.66 67.45
-Router 16.47 4.11 43.04 22.69 10.92 54.55
-25 Skip Decode Router + LoRA 16.85 7.62 32.33 27.81 15.74 42.01
-Router 9.14 3.51 27.64 7.01 3.41 29.51
-Random Skip Router + LoRA 8.95 1.07 30.97 25.12 9.85 45.53
-Router 5.32 0.69 27.22 10.07 3.69 31.50
-Unified Skip Router + LoRA 15.14 2.65 39.81 30.56 12.27 42.30
-Router 10.21 1.62 30.86 15.21 5.51 31.63
-FiRST (Ours) Router + LoRA 18.89 4.06 45.38 32.92 13.74 41.66
-Router 9.51 1.44 29.08 10.46 3.55 27.83
-Table 2: Machine Translation Results for LLaMA-3.2-3B: BLEU-1, BLEU-2, and COMET scores are shown for English-to-German and English-to-Chinese tasks at different skip levels.
-FiRST consistently achieves the highest BLEU-1 and COMET scores for English-to-German at both skipping percentages and highest BLEU-1 for English-to-Chinese.
-Skip (%) Model Type EM F1
-0 Original Model Base + LoRA 73.93 85.99
-Base 19.46 36.73
-10 Skip Decode R + LoRA 60.14 65.33
-Router 16.38 31.48
-Random Skip R + LoRA 65.73 80.08
-Router 18.25 33.75
-Unified Skip R + LoRA 55.54 74.58
-Router 17.39 32.91
-FiRST (Ours) R + LoRA 70.85 83.61
-Router 14.58 31.52
-20 Skip Decode R + LoRA 45.00 55.10
-Router 10.68 26.69
-Random Skip R + LoRA 47.79 66.37
-Router 6.71 22.46
-Unified Skip R + LoRA 52.87 69.28
-Router 18.18 32.51
-FiRST (Ours) R + LoRA 60.60 75.49
-Router 13.21 27.48
-Skip (%) Model Type EM F1
-0 Original Model wLoRA 73.07 84.17
-Base 18.92 37.74
-10 Skip Decode R + LoRA 60.79 75.00
-Router 20.00 31.55
-Random Skip R + LoRA 64.78 77.27
-Router 13.76 28.59
-Unified Skip R + LoRA 65.03 77.53
-Router 13.16 32.31
-FiRST R + LoRA 69.44 81.35
-Router 12.79 28.37
-20 Skip Decode R + LoRA 40.12 40.00
-Router 20.45 37.62
-Random Skip R + LoRA 11.32 38.34
-Router 6.75 15.51
-Unified Skip R + LoRA 37.39 52.49
-Router 7.81 18.20
-FiRST R + LoRA 39.70 54.59
-Router 5.52 15.33
-Table 3: Quality Analysis on Question Answering (SQuAD v1.1): LLaMA-3-8B (left) and LLaMA-3.2-3B (right).
-Exact Match (EM) and F1 scores are reported for varying skipping levels.
-Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
+Skip (%) Model Type SQuAD NQ
+EM F1 EM
+LLaMA-3-8B
+0 Original Model Base + LoRA 73.93 85.99 51.88
+Base 19.46 36.73 37.40
+10 Skip Decode R + LoRA 60.14 65.33 41.98
+Router 16.38 31.48 33.22
+Random Skip R + LoRA 65.73 80.08 44.95
+Router 18.25 33.75 34.09
+Unified Skip R + LoRA 55.54 74.58 45.91
+Router 17.39 32.91 33.64
+FiRST (Ours) R + LoRA 70.85 83.61 47.85
+Router 14.58 31.52 33.53
+20 Skip Decode R + LoRA 45.00 55.10 26.62
+Router 10.68 26.69 14.39
+Random Skip R + LoRA 47.79 66.37 28.96
+Router 6.71 22.46 27.40
+Unified Skip R + LoRA 52.87 69.28 25.30
+Router 18.18 32.51 23.57
+FiRST (Ours) R + LoRA 60.60 75.49 32.22
+Router 13.21 27.48 18.02
+LLaMA-3.2-3B
+0 Original Model Base + LoRA 73.07 84.17 40.50
+Base 18.92 37.74 30.10
+10 Skip Decode R + LoRA 60.79 75.00 31.74
+Router 20.00 31.55 21.51
+Random Skip R + LoRA 64.78 77.27 36.02
+Router 13.76 28.59 22.76
+Unified Skip R + LoRA 65.03 77.53 32.90
+Router 13.16 32.31 21.03
+FiRST (Ours) R + LoRA 69.44 81.35 37.82
+Router 12.79 28.37 22.55
+20 Skip Decode R + LoRA 40.12 40.00 27.28
+Router 20.45 37.62 14.73
+Random Skip R + LoRA 11.32 38.34 26.60
+Router 6.75 15.51 13.24
+Unified Skip R + LoRA 37.39 52.49 26.84
+Router 7.81 18.20 16.09
+FiRST (Ours) R + LoRA 39.70 54.59 29.87
+Router 5.52 15.33 17.51
+Table 2: Quality Analysis on Question Answering tasks using LLaMA-3-8B and LLaMA-3.2-3B across SQuAD (Exact Match and F1) and Natural Questions (Exact Match).
+Results are shown for multiple skip strategies and levels.
+FiRST (Ours) consistently performs best under skipping, demonstrating strong robustness.
+R + LoRA indicates Router Augmentation followed by LoRA fine-tuning.
 Skip (%) Model Type BERT R-1 R-L
-0 Original Model wLoRA 84.87 28.46 16.99
+LLaMA-3-8B
+0 Original Model Base + LoRA 84.87 28.46 16.99
 Base 82.29 23.49 14.66
@@
 Router 81.25 20.2 13.01
-27 Skip Decode R + LoRA 79.92 10.67 10.32
-Router 77.27 9.59 7.00
-Random Skip R + LoRA 76.40 11.45 7.89
-Router 77.45 12.56 9.08
-Unified Skip R + LoRA 80.28 15.94 9.89
-Router 77.43 10.97 7.68
-FiRST (Ours) R + LoRA 77.5 14.65 10.45
-Router 75.6 9.39 6.92
+20 Skip Decode R + LoRA 82.57 20.41 14.87
+Router 81.62 13.48 9.19
+Random Skip R + LoRA 81.39 21.57 13.83
+Router 79.23 15.51 10.93
+Unified Skip R + LoRA 82.93 22.3 13.37
+Router 80.32 16.51 11.15
+FiRST (Ours) R + LoRA 82.8 27.65 17.84
+Router 79.32 16.28 10.85
 Skip (%) Model Type BERT R-1 R-L
-0 Original Model wLoRA 84.89 28.37 17.02
+LLaMA-3.2-3B
+0 Original Model Base + LoRA 84.89 28.37 17.02
 Base 71.85 19.34 12.00
@@
 Table 3: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels.
-Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
-We conduct experiments on three benchmark tasks namely Machine Translation, Text Summarization and Question Answering demonstrating both robustness and scalability of FiRST.
+Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework).
+We conduct experiments on three benchmark tasks—Machine Translation, Text Summarization, and Question Answering—demonstrating the robustness and scalability of FiRST across diverse settings.
+We base our task selection on prior work in the field to ensure a fair and meaningful comparison Liu et al.
+(2024); Del Corro et al.
+(2023); Schuster et al.
+(2022b).
 Datasets: For machine translation, we use WMT development sets (2017–2020) for English-to-Chinese and English-to-German tasks, evaluating performance on the WMT 2022 test set, which covers diverse domains such as news, social media, e-commerce, and conversational contexts.
-For summarization, we use the CNN/DailyMail dataset, with 4000 randomly selected training samples and evaluation on the standard test set of 11,490 samples.
-In Question Answering, we utilize SQuAD v1.0, training on 4000 randomly selected samples from the 100k question-answer pairs and evaluating on the validation set, as test set labels are unavailable.
+For summarization, we use the CNN/DailyMail dataset, with 4,000 randomly selected training samples and evaluation on the standard test set of 11,490 samples.
+For question answering (QA), we utilize SQuAD v1.1 and Natural Questions (NQ).
+We train on 4,000 randomly selected samples from each dataset and evaluate on their respective validation sets as test set labels are unavailable.
+For NQ, we incorporate a retrieval step before answer generation, retrieving relevant passages as context.
 Appendix A.2 contains detailed descriptions.
-Evaluation Metrics: We use standard metrics to evaluate the quality of generated output in each of the three tasks.
-For Machine Translation, we benchmark using BLEU scores and COMET where the latter provides a more nuanced assessment beyond $n$-gram used in BLEU.
-For Summarization, we use ROUGE scores and BERT Score - as before, the latter captures meaning-based similarity.
-For Question Answering, we use Exact Match and F1 Score as the metrics to benchmark the output quality.
-Finally, for benchmarking latency, we look at the TPOT (Time Per Output Token).
-TPOT evaluates the average time taken to produce each output token and is calculated for GPU to gauge overall decoding performance.
-Appendix A.3 contains a detailed description of evaluation metrics.
-Additionally, hyper-parameters used during training and inference can be found in Appendix A.4.
+Evaluation Metrics: We employ standard metrics to assess output quality across three distinct tasks.
+For Machine Translation, we benchmark performance using BLEU and COMET; COMET is included for its more nuanced assessment capabilities beyond the n-gram overlap measured by BLEU.
+Summarization quality is evaluated with ROUGE scores and BERT Score, the latter capturing meaning-based similarity.
+For Question Answering, we use Exact Match (EM) and F1 Score on the SQuAD dataset, and report Exact Match (EM) for Natural Questions (NQ).
+Finally, to benchmark latency, we measure Time Per Output Token (TPOT) on GPU, which gauges overall decoding performance.
+Detailed descriptions of all evaluation metrics are available in Appendix A.3.
+Hyperparameters used during training and inference are provided in Appendix A.4.
 5.1 Baselines for comparison
@@
 (2023) method that features a monotonic decrease in processing layers, enabling later tokens to leverage the computational resources used for earlier ones.
-  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method that relies on using a heuristic-based strategy for retaining layers at fixed intervals.
+  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method relies on using a heuristic-based strategy for retaining layers at fixed intervals.
 We replicate the algorithm in (Liu et al., 2024) and compare performance both with and without LoRA fine-tuning across various skipping percentages.
 5.2 Experimental Results on Different Tasks
-WMT: Our experiments demonstrate that FiRST performs consistently well in different machine translation tasks (EN-to-DE, EN-to-ZH) across different models.
-For LLaMA-3.8B, for $15\%$ skipping, FiRST achieves a latency improvement of 10-12% on TPOT (see Tables 1 and 5) while being competitive to the base model + LoRA (gold reference) in quality.
-In most cases, it significantly outperforms other layer skipping strategies (Random, Skip Decode, Unified Skipping) and in other cases, it is comparable in quality.
-In comparison to the gold output (Base + LoRA), FiRST is $\geq 80\%$ in COMET scores (semantic metric), $\geq 85\%$ in BLEU-1 scores, and $\geq 75\%$ in BLEU-2 scores (syntax metrics) for both EN-to-DE and EN-to-ZH translations.
+WMT: For LLaMA-3-8B, at $15\%$ skipping, FiRST achieves a latency improvement of upto 10% on TPOT (see Tables 1 and 4).In most cases, it significantly outperforms other layer skipping strategies (Skip Decode, Random and Unified Skipping) and in other cases, it is comparable in quality.
+When compared to the gold output (Base + LoRA), FiRST generally retains a high percentage of the quality, for instance, often achieving $\geq 80\%$ of the COMET score for LLaMA-3-8B, and $\geq 70\%$ in BLEU scores.
 For $25\%$ skipping, FiRST achieves significant improvement in quality over other strategies, in almost all metrics, while achieving $\sim 18\%$ reduction in TPOT.
-For LLaMA-3.2-3B, the latency improvement is $\sim 10\%$, with quality scores being significantly higher than other layer skipping strategies (Table 6).
-It is within $65-85\%$ of BLEU-1 and COMET scores (Table 2) for EN-DE and EN-ZH of the gold (LoRA fine-tuned base model).
-CNN/DailyMail Dataset: For LLaMA-3-8B, at roughly 15% skipping level, our method outperforms the base model + LoRA (Table: 4) while obtaining a 12% improvement in TPOT (Table 5).
+For LLaMA-3.2-3B, for similar latency improvement ($\sim 10\%$), the COMET scores are significantly higher that other baselines while the BLEU scores are comparable.
+(Table 4).
+It remains within $65-85\%$ of BLEU and COMET scores (Table 1) achieved by the gold standard.
+CNN/DailyMail Dataset: For LLaMA-3-8B, at $\sim 15\%$ skipping, our method outperforms the Base + LoRA setting(Table: 3) while obtaining a 12% improvement in TPOT (Table 4).
 For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT.
 At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
-Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
-TPOT TPOT
-Base + LoRA 0 1x 1x
-R + LoRA 15 0.90x 0.88x
-R + LoRA 25 0.82x 0.83x
-R + LoRA 35 0.69x 0.68x
-[…]
-TPOT
-Base + LoRA 0 1x
-R + LoRA 15 0.88x
-R + LoRA 20 0.81x
-R + LoRA 27 0.76x
-Table 5: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
-Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
-TPOT TPOT
-Base + LoRA 0 1x 1x
-R + LoRA 15 0.90x 0.91x
-R + LoRA 25 0.78x 0.75x
-R + LoRA 35 0.69x 0.74x
-[…]
-TPOT
-Base + LoRA 0 1x
-R + LoRA 15 0.88x
-R + LoRA 24 0.79x
-R + LoRA 28 0.77x
-Table 6: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
-[…]
-TPOT
-Base + LoRA 0 1x
-R + LoRA 15 0.94x
-R + LoRA 24 0.83x
-R + LoRA 28 0.74x
-[…]
-TPOT
-Base + LoRA 0 1x
-R + LoRA 15 0.94x
-R + LoRA 24 0.84x
-R + LoRA 28 0.72x
-Table 7: TPOT variation of LLaMA-3-8B (left) and LLaMA-3.2-3B (right) on SQuAD dataset for FiRST.
-These values are relative to the LoRA fine-tuned base model.
+Model size Model Type WMT CNN/DM SQuAD Natural Questions
+Skip (%) Eng→De Eng→Zh Skip (%) TPOT Skip (%) TPOT Skip (%) TPOT
+LLaMA-3–8.B Base + LoRA 0 0 0 0
+R + LoRA 15 0.90× 0.88× 15 0.88× 10 0.95× 10 0.96×
+R + LoRA 25 0.82× 0.83× 20 0.81× 20 0.78× 20 0.80×
+LLaMA-3.2–3B Base + LoRA 0 0 0 0
+R + LoRA 15 0.90× 0.91× 15 0.88× 15 0.94× 15 0.94×
+R + LoRA 25 0.78× 0.75× 24 0.79× 24 0.83× 24 0.84×
+Table 4: TPOT variation across all datasets for FiRST.
+The reported values are relative to the LoRA fine-tuned base model.
+R + LoRA indicates Router Augmentation followed by LoRA fine-tuning.
 Fine-tuning improves TPOT and quality significantly.
-SQuAD Dataset: For the LLaMA-3-8B model, FiRST is $>95\%$ in output quality of gold (base + LoRA) (Table 3), with overall latency gains of 6-16% (Table 7).
+SQuAD Dataset: For the LLaMA-3-8B model, FiRST(at 10% skip level) maintains over 95% of the performance of the gold standard (Base + LoRA without skipping)(Table 2), with overall latency gains upto 16% (Table 4).
 It is significantly better in quality than all other baselines across all metrics for different levels of skipping.
-For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 3) with gains in latency of 6-16% overall (Table 7) over the LoRA fine-tuned base model.
+For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model.
 Moreover, it is better than all other layer skipping strategies across all metrics.
+Natural Questions Dataset: For the LLaMA-3-8B model, FiRST retains over $92\%$ of the EM achieved by the non-skipping gold standard (Table 2), with overall latency gains of 5-12% (Table 4).
+It is significantly better in quality than all other baselines for different levels of skipping.
+For LLaMA-3.2-3B, again FiRST is $>93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model.
+Moreover, it is better than all other layer skipping strategies.
+While more sophisticated retrieval strategies could further enhance performance, our goal was to demonstrate that FiRST maintains quality performance and latency gains even in a retrieval-based QA setting.
 Detailed results for an additional skipping percentage are provided in Appendix A.7.
@@
 For LLaMA-3-8B at a 15% skipping rate, layers 7–9 and 21 are fully skipped in English-to-German, with partial skipping in layer 18.
-In English-to-Chinese, layers 7–9, 16, and 21 are fully skipped, while layer 20 is partially skipped.
-For Summarization, layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19, 21, and 26.
-Some layers are skipped less than 10%, indicating their necessity for specific sequences.
-This task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped, and skipping patterns depend on the input.
+Figure 3 for summarization shows that only some layers in the middle of the network, specifically layers 19 to 27, are skipped.
+The early layers and the last few layers (28–32) are never skipped.
+Layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19 and 21.
+Some layers are skipped less than 10% of times, indicating their necessity for specific sequences.
+This shows that it’s important to learn which layers to skip based on each input, as the skipping pattern is not the same for every sample.
+The task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped on SQuAD while layers 12, 23 are fully skipped in case of NQ dataset, and skipping patterns depend on the input.
 Detailed statistics are in Appendix A.6.
-Computational overhead of Routers: We also examine the computational overhead introduced by training the routers.
-This overhead remains minimal, as the router parameters are significantly smaller than the total number of trainable parameters in each model (0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B).
-The proportion of the total computational time spent on the router operations compared to the entire forward pass of the model is close to 0.3% for both models suggesting their overhead does not significantly affect the efficiency of the model during inference or training.
+Furthermore, the router is trained in a task-aware manner, ensuring that skipping decisions align with task complexity.
+We also conduct experiments to validate the reusability and generalizability of routers (see Appendix: A.8) across different datasets for the same underlying task.
+Figure 3: LLaMA-3-8B skipping statistics at 15% skipping rate on summarization task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
+Computational overhead of Routers: We aim to improve the efficiency of pre-trained models by introducing routers, lightweight linear classifiers that help decide whether to skip certain layers during inference.
+A common concern with adding such components is the potential increase in computational cost.
+However, our analysis shows that the additional parameters and operations introduced by the routers account for a very small fraction of the total model computation, only 0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B in terms of parameter count, and around 0.3% of the total forward pass time.
+This fixed cost remains negligible compared to the overall model execution time.
+As a result, the overall latency gain primarily depends on how much computation is skipped, rather than the routing mechanism itself.
+Given similar levels of skipping, different methods tend to show comparable efficiency gains, so the key focus becomes minimizing the performance drop relative to the original model.
+Our results show that the proposed routing mechanism achieves this balance effectively, preserving latency improvements while maintaining model quality.
 6 Conclusion
-We propose a new framework FiRST for layer selection corresponding to input sequence and task towards reducing latency in a quality aware manner.
-This is sequence dependent and operates in a KV cache compatible manner.
-For an optimal skipping rate of around 15%, FiRST achieves 10-20% reduction in latency while being quality neutral (approximately 80% or more in quality metrics compared to base model) on multiple tasks/model architectures such as Machine Translation, Summarization and Question Answering, on well known open source datasets, and some cases, even improving upon base model, while significantly outperfoming other layer selection strategies on most quality metrics.
+We propose a new framework, FiRST, for layer selection that adapts to the input sequence and task, aiming to reduce latency in a quality-aware manner.
+This approach is sequence-dependent and operates compatibly with KV caching.
+With an optimal layer skipping rate of around 15%, FiRST achieves a 10-20% reduction in latency.
+This speedup is achieved while remaining quality-neutral, maintaining approximately 80% or more of the base model’s performance on quality metrics across multiple tasks (Machine Translation, Summarization, Question Answering) and model architectures on well-known open-source datasets.
+Furthermore, our method significantly outperforms other layer selection strategies on most quality metrics.
 7 Limitations
@@
 Early Exit:
-Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022) where adaptive compute is used for different parts of the token sequence.
+Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022a; Wu et al., 2020) where adaptive compute is used for different parts of the token sequence.
 While these approaches have been popular for encoder-only models which processes the entire sequence as a whole, they have faced challenges in generation tasks.
@@
 (2023) have imposed (query aware/ query agnostic) sparsity in the KV cache matrices to speed up self-attention mechanism via clever selection of the critical tokens necessary from the KV cache.
+Mixture of Experts
+Mixture of Experts (MoE) is a well-established technique for improving the efficiency and capacity of deep learning models by conditionally activating subsets of parameters for different inputs.
+MoE-based transformer models, such as Switch Transformers (Fedus et al., 2021) and GShard (Lepikhin et al., 2020), employ a gating mechanism to route tokens to a subset of expert layers, thereby significantly reducing computational costs while maintaining expressivity.
+These architectures are designed to scale up model capacity without a proportional increase in inference cost.
+The main goal is to scale up the parameters while maintaining the cost of pre-training and inference.
+Although MoE and input-adaptive layer skipping share the goal of selective computation, they differ in fundamental ways.
+MoE dynamically routes tokens to different experts at the layer level, whereas layer skipping focuses on bypassing entire layers in the transformer stack based on the input query.
+MoE models typically maintain a full-depth model structure, leveraging sparse activation to reduce computational overhead, whereas layer-skipping methods explicitly modify the depth of computation for different inputs.
+Simply put, the techniques are orthogonal - layer skipping can be used in tandem with Moe to further reduce the depth of computation wherever possible at an expert level
 A.2 Details of Datasets
+WMT Summarization Question Answering
+[…]
+Train 3,505 8,983 3,400 3,400 3,400
+Validation 876 998 600 600 600
+Test 2,038 2,038 11,490 10,570 7,830
+Table 5: Train, validation, and test splits for machine translation (WMT), summarization (CNN/DM), and question answering (SQuAD, NQ) tasks.
 Machine Translation: For translation tasks, namely English-to-Chinese and English-to-German, we employ the WMT development sets from 2017 to 2020 for training/fine-tuning following the methodology outlined in previous studies (Liu et al., 2023a; Jiao et al., 2023).
@@
 These domains include news, social media, e-commerce, and conversational contexts.
-(Details in Appendix: A.5, Table: 8).
 Summarization: We use the popular CNN-DailyMail (CNN/DM) (Hermann et al., 2015) dataset which is a large collection (over 300k) of text summarization pairs, created from CNN and Daily Mail news articles.
@@
 During training in our framework, the number of trainable parameters is small in both phases - therefore a small subset of data points is sufficient for training.
-Inference is performed on the standard test set with 11,490 samples.
-Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.0) (Rajpurkar et al., 2016), a widely-used benchmark for machine Question Answering.
+Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.1) (Rajpurkar et al., 2016), a widely-used benchmark for Machine Question Answering.
 The dataset consists of over 100k question-answer pairs posed by crowd-workers on a set of over 500 Wikipedia articles.
-Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage) and the corresponding answer (a text span from the corresponding reading passage).
-Similarly to the CNN/DM dataset, 4k samples are chosen as random to train both routers and LoRA.
-The training and validation split contains 87,599 and 10,570 samples respectively.
+Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage), and the corresponding answer(a text span from the corresponding reading passage).
+Similarly to the CNN/DM dataset, 4k samples are chosen at random to train both routers and LoRA.
+The training and validation splits contain 87,599 and 10,570 samples, respectively.
 Evaluation is performed on the validation set Schuster et al.
-(2022) as the test set labels are not publicly released.
+(2022a) as the test set labels are not publicly released.
+Natural Questions (NQ): We use the Natural Questions dataset (Kwiatkowski et al., 2019), which consists of real queries issued to Google Search paired with relevant Wikipedia articles.
+Each example in NQ contains a query (an actual user question), a context(Wikipedia article), and two types of answers: a long answer (typically a paragraph) and a short answer (a specific text span, when available).
+We implement a naïve RAG solution using a simple S-BERT model to generate embeddings for the query and passages and retrieve the top 5 most relevant passages based on similarity.
+Once retrieved, these passages are used as context to answer the query and then compared against gold answers.
+Similar to SQuAD, 4k samples are randomly chosen to train both routers and LoRA.
+The training set contains over 300k examples, with evaluation performed on 7.83k validation samples containing short answers.
+For further details on training-testing split, refer to table: 5
 A.3 Evaluation Metrics
@@
 BLEU compares n-grams of the candidate translation to n-grams of the reference translation, providing a score between 0 and 1, with higher scores indicating better translations.
-In this evaluation, NLTK BLEU is employed, focusing on BLEU-1 and BLEU-2 scores.
+In this evaluation, NLTK BLEU is employed and We report BLEU-1, BLEU-2, and the cumulative BLEU score, which is computed as the geometric mean of individual n-gram precision scores from unigram to 4-gram.
+To mitigate the issue of zero counts for higher-order n-grams, we apply the smoothing strategy utilized in Post (2018).
   • COMET: COMET (Cross-lingual Optimized Metric for Evaluation of Translation) is used to assess translation quality further.
@@
 A.4 Training and Inference Setup
-  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively.
-Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization.
-The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps.
-The learning rate is set between $1e^{-4}$ and $3e^{-4}$ a cosine scheduler is used to adjust the learning rate.
-Gradients are accumulated after 5 steps and the regularization coefficient $\lambda$ is fixed at 0.01.
-For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor (lora_alpha) of 32.
-We set the non-skipping penalization regularizer hyper-parameter $\beta=\alpha/3$ after tuning to maintain skipping level - where $\alpha$ is the non-skipping penalization hyperparameter used in the first phase and set according to the target speedup ratio.
-For translation, the maximum sequence length is set to 128 for router training and 256 for LoRA training.
-Similarly, for summarization, the maximum sequence length is set to 500 and 700 respectively.
-For Question Answering, this length is set to 512.
-Prompts for the different tasks regarding training/inference are shown in Appendix A.6.
+[…]
+  • Joint training considerations: The decision to split training into two phases—first training the router and then the LoRA modules—was made to ensure better stability during optimization.
+In our experiments, this approach allowed the router to establish a robust skipping strategy without interference from the fine-tuning of LoRA modules, which otherwise introduces instability.
+While training both the router and LoRA modules together is theoretically possible and could eliminate the need for reapplying the non-skip penalty in the second phase, initial experiments revealed that it led to suboptimal convergence and conflicting gradients.
+This was particularly evident when the router’s decisions were not yet stable, causing inconsistencies in the joint optimization process.
   • Inference settings: For all the tasks, we set the temperature to 0.8 and enable top-k sampling over 10 tokens.
-The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD.
+The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD and NQ.
 Caching is turned on during inference.
-A.5 Training and testing split
-WMT Summarization RC
-Eng-to-German Eng-to-Chinese CNN/DM SQuAD
-Train 3505 8983 3400 3400
-Validation 876 998 600 600
-Test 2038 2038 11490 10570
-Table 8: Train-Validation-Test split for WMT, CNN/DM and SQuAD datasets
 Figure 4: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on English-to-German Machine Translation Task.
@@
 Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
-A.6 Prompt Details
+Figure 8: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 10% skipping rate on Natural Questions Answering Task.
+Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
+A.5 Prompt Details
 The prompt structures used for both training and inference are as follows:
@@
     ### Response:
-  • For the Question Answering task (used in SQuAD dataset), the following prompt structure is utilized:
+  • For the Question Answering task (used in SQuAD/NQ dataset), the following prompt structure is utilized:
     ### Instruction:
@@
 […]
-A.7 Layer-wise Skipping Statistics
-Table 15 present the fraction of sequences that skip a particular block during the task for the LLaMA-3-8B model.
-If the corresponding cell in a row shows a value of 80.00, it implies that 80% of the sequences skip this block.
-It is important to note that the decision regarding which block to skip varies across different datasets and tasks.
-Additionally, partial skipping in some blocks, with varying percentages, suggests that while some sequences consider the layer important, others do not and therefore skip it during the decoding phase.
-Figures 3, 4, 5, and 6 depict the distribution of skipped blocks when the model is configured to skip approximately 15% of the layers.
-These plots demonstrate the variation in block skipping based on the task and dataset, highlighting the differences in block importance.
-The bar plots further show that the determination of important layers depends not only on the dataset but also on the model architecture.
-For instance, Figure 3 illustrates that layers 7 and 9 are entirely skipped in LLaMA-3-8b, while they are rarely skipped in LLaMA-3.2-3b.
-A.8 Detailed Result Table
-Tables 9 and 11 present the detailed results for the LLaMA-3.2-3B model, while Tables 10 and 12 summarize the performance of the LLaMA-3-8B model for an additional skipping percentage on Machine Translation Task.
-The results are reported using BLEU (BLEU-1, BLEU-2) and COMET metrics, highlighting performance across different skipping percentages.
-Similarly, Table 13 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L.
-Lastly, Table 14 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
-Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 37.57 17.43 89.72
-Base 31.19 13.67 81.66
-15 Skip Decode Router + LoRA 23.24 11.07 44.58
-Router 9.93 3.89 32.74
-Random Skip Router + LoRA 18.99 4.78 47.26
-Router 12.11 2.77 36.30
-Unified Skip Router + LoRA 21.65 5.93 44.72
-Router 16.41 3.99 39.81
-FiRST (Ours) Router + LoRA 24.68 7.13 60.29
-Router 16.47 4.11 43.04
-25 Skip Decode Router + LoRA 16.85 7.62 32.33
-Router 9.14 3.51 27.64
-Random Skip Router + LoRA 8.95 1.07 30.97
-Router 5.32 0.69 27.22
-Unified Skip Router + LoRA 15.14 2.65 39.81
-Router 10.21 1.62 30.86
-FiRST (Ours) Router + LoRA 18.89 4.06 45.38
-Router 9.51 1.44 29.08
-35 Skip Decode Router + LoRA 1.62 0.36 23.30
-Router 1.71 0.27 19.34
-Random Skip Router + LoRA 4.72 0.15 25.08
-Router 1.96 0.06 21.18
-Unified Skip Router + LoRA 1.18 0.05 19.65
-Router 0.90 0.03 20.36
-FiRST (Ours) Router + LoRA 9.47 1.29 27.45
-Router 5.74 0.56 25.03
-Table 9: Machine Translation Results for English to German on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
-Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 41.78 21.74 93
-Base 37.17 18.57 87.13
-15 Skip Decode Router + LoRA 23.04 10.52 55.62
-Router 3.99 1.18 23.33
-Random Skip Router + LoRA 30.43 10.98 66.25
-Router 26.54 8.77 60.27
-Unified Skip Router + LoRA 28.92 10.64 59.34
-Router 23.23 7.85 59.26
-FiRST (Ours) Router + LoRA 38.01 17.89 82.14
-Router 28.83 11.8 67.74
-25 Skip Decode Router + LoRA 13.67 6.00 31.47
-Router 3.24 0.92 21.55
-Random Skip Router + LoRA 6.01 0.91 29.71
-Router 3.65 0.49 29.95
-Unified Skip Router + LoRA 15.67 3.36 31.69
-Router 12.58 2.65 32.15
-FiRST (Ours) Router + LoRA 17.84 4.14 34.95
-Router 9.67 1.37 26.01
-35 Skip Decode Router + LoRA 5.55 0.33 23.85
-Router 3.03 0.82 20.03
-Random Skip Router + LoRA 1.80 0.12 25.56
-Router 1.44 0.06 25.34
-Unified Skip Router + LoRA 6.44 0.77 22.05
-Router 3.92 0.51 22.88
-FiRST (Ours) Router + LoRA 6.39 0.42 19.96
-Router 3.7 0.14 21.41
-Table 10: Machine Translation Results for English to German on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
-Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 51.81 30.04 79.13
-Base 32.10 17.92 61.84
-15 Skip Decode Router + LoRA 38.14 21.68 46.70
-Router 9.59 4.84 34.14
-Random Skip Router + LoRA 38.75 17.39 57.79
-Router 13.56 5.64 35.86
-Unified Skip Router + LoRA 36.96 17.35 57.10
-Router 22.43 9.37 45.16
-FiRST (Ours) Router + LoRA 45.66 23.66 67.45
-Router 22.69 10.92 54.55
-25 Skip Decode Router + LoRA 27.81 15.74 42.01
-Router 7.01 3.41 29.51
-Random Skip Router + LoRA 25.12 9.85 45.53
-Router 10.07 3.69 31.50
-Unified Skip Router + LoRA 30.56 12.27 42.30
-Router 15.21 5.51 31.63
-FiRST (Ours) Router + LoRA 32.92 13.74 41.66
-Router 10.46 3.55 27.83
-35 Skip Decode Router + LoRA 2.84 1.10 22.30
-Router 2.36 1.04 18.32
-Random Skip Router + LoRA 4.17 1.52 32.81
-Router 2.11 0.76 23.63
-Unified Skip Router + LoRA 3.26 0.52 19.12
-Router 2.20 0.37 19.82
-FiRST (Ours) Router + LoRA 20.01 6.96 28.10
-Router 6.54 1.78 23.13
-Table 11: Machine Translation Results for English to Chinese on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
-Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
-0 Original Model Base + LoRA 56.94 35.56 82.66
-Base 38.02 22.46 68.95
-15 Skip Decode Router + LoRA 28.73 15.84 55.98
-Router 4.74 2.33 21.75
-Random Skip Router + LoRA 47.88 25.75 67.32
-Router 36.60 18.66 59.89
-Unified Skip Router + LoRA 46.61 25.01 69.58
-Router 27.28 13.35 54.57
-FiRST (Ours) Router + LoRA 48.35 26.57 68.63
-Router 17.55 8.68 42.76
-25 Skip Decode Router + LoRA 20.03 10.85 33.85
-Router 3.78 1.84 20.93
-Random Skip Router + LoRA 11.69 4.66 27.73
-Router 7.37 2.81 35.16
-Unified Skip Router + LoRA 34.90 15.75 50.59
-Router 17.74 7.35 38.74
-FiRST (Ours) Router + LoRA 35.79 15.66 56.92
-Router 11.01 3.23 25.45
-35 Skip Decode Router + LoRA 12.24 6.27 25.23
-Router 3.64 1.74 22.84
-Random Skip Router + LoRA 7.38 2.19 27.35
-Router 4.47 1.12 29.46
-Unified Skip Router + LoRA 7.51 2.10 20.25
-Router 3.87 1.06 21.24
-FiRST (Ours) Router + LoRA 15.66 3.95 26.80
-Router 6.13 1.54 22.89
-Table 12: Machine Translation Results for English to Chinese on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
+A.6 Layer-wise Skipping Statistics
+To illustrate broader layer-wise behaviors, Figures 4 through 8 show block-wise skip rates observed when the models were configured for an average skip rate of approximately 15%.
+These figures highlight how individual layers exhibit distinct skipping behaviors across various tasks, including translation (English-German, English-Chinese), summarization (CNN/DM), and question answering (SQuAD, Natural Questions).
+It also reveals that tasks with similar processing requirements (e.g., different translation language pairs) exhibit comparable skipping patterns.
+Providing a more granular view for specific models, Table 8 reports the fraction of sequences that skip each block under LoRA adaptation for the LLaMA-3-8B and LLaMA-3.2-3B models.
+These fractions are presented as decimals, so a value of 0.10 signifies that 10% of sequences skip the corresponding block.
+Skip (%) Model Type Config BLEU-1 BLEU-2 BLEU COMET
+LLaMA-3.2-3B
+0 Original Model Base + LoRA 0.376 0.174 0.160 89.72
+Base 0.312 0.137 0.125 81.66
+15 Skip Decode Router + LoRA 0.207 0.069 0.105 44.58
+Router 0.099 0.039 0.055 32.74
+Random Skip Router + LoRA 0.190 0.048 0.079 47.26
+Router 0.121 0.028 0.048 36.30
+Unified Skip Router + LoRA 0.216 0.059 0.095 44.72
+Router 0.164 0.040 0.067 39.81
+FiRST (Ours) Router + LoRA 0.247 0.071 0.113 60.29
+Router 0.165 0.041 0.069 43.04
+25 Skip Decode Router + LoRA 0.168 0.047 0.077 32.33
+Router 0.091 0.035 0.050 27.64
+Random Skip Router + LoRA 0.090 0.011 0.023 30.97
+Router 0.053 0.007 0.015 27.22
+Unified Skip Router + LoRA 0.151 0.027 0.050 39.81
+Router 0.102 0.016 0.032 30.86
+FiRST (Ours) Router + LoRA 0.189 0.041 0.072 45.38
+Router 0.095 0.014 0.029 29.08
+35 Skip Decode Router + LoRA 0.058 0.011 0.018 23.30
+Router 0.017 0.003 0.005 19.34
+Random Skip Router + LoRA 0.047 0.001 0.007 25.08
+Router 0.020 0.001 0.003 21.18
+Unified Skip Router + LoRA 0.012 0.000 0.002 19.65
+Router 0.009 0.000 0.001 20.36
+FiRST (Ours) Router + LoRA 0.095 0.013 0.027 27.45
+Router 0.057 0.006 0.014 25.03
+LLaMA-3-8B
+0 Original Model Base + LoRA 0.418 0.217 0.199 93.00
+Base 0.372 0.186 0.169 87.13
+15 Skip Decode Router + LoRA 0.230 0.105 0.094 55.62
+Router 0.040 0.012 0.019 23.33
+Random Skip Router + LoRA 0.304 0.110 0.097 66.25
+Router 0.265 0.088 0.132 60.27
+Unified Skip Router + LoRA 0.289 0.106 0.153 59.34
+Router 0.232 0.079 0.117 59.26
+FiRST (Ours) Router + LoRA 0.380 0.179 0.161 82.14
+Router 0.288 0.118 0.108 67.74
+25 Skip Decode Router + LoRA 0.118 0.052 0.070 31.47
+Router 0.032 0.009 0.015 21.55
+Random Skip Router + LoRA 0.060 0.009 0.018 29.71
+Router 0.037 0.005 0.011 29.95
+Unified Skip Router + LoRA 0.157 0.034 0.060 31.69
+Router 0.126 0.027 0.048 32.15
+FiRST (Ours) Router + LoRA 0.178 0.041 0.071 34.95
+Router 0.097 0.014 0.029 26.01
+35 Skip Decode Router + LoRA 0.049 0.020 0.028 23.85
+Router 0.030 0.008 0.013 20.03
+Random Skip Router + LoRA 0.018 0.001 0.004 25.56
+Router 0.014 0.001 0.002 25.34
+Unified Skip Router + LoRA 0.064 0.008 0.017 22.05
+Router 0.039 0.005 0.011 22.88
+FiRST (Ours) Router + LoRA 0.064 0.004 0.012 19.96
+Router 0.037 0.001 0.005 21.41
+Table 6: Machine Translation Results for English to German on LLaMA-3.2-3B and LLaMA-3-8B: BLEU and COMET scores for various skipping strategies.
+Skip (%) Model Type Config BLEU-1 BLEU-2 BLEU COMET
+LLaMA-3.2-3B
+0 Original Model Base + LoRA 0.518 0.300 0.278 79.13
+Base 0.321 0.179 0.166 61.84
+15 Skip Decode Router + LoRA 0.381 0.217 0.200 46.70
+Router 0.096 0.048 0.062 34.14
+Random Skip Router + LoRA 0.387 0.174 0.154 57.79
+Router 0.136 0.056 0.077 35.86
+Unified Skip Router + LoRA 0.370 0.173 0.157 57.10
+Router 0.224 0.094 0.087 45.16
+FiRST (Ours) Router + LoRA 0.457 0.237 0.218 67.45
+Router 0.227 0.109 0.100 54.55
+25 Skip Decode Router + LoRA 0.278 0.157 0.147 42.01
+Router 0.070 0.034 0.045 29.51
+Random Skip Router + LoRA 0.251 0.099 0.137 45.53
+Router 0.101 0.037 0.053 31.50
+Unified Skip Router + LoRA 0.306 0.123 0.108 42.30
+Router 0.152 0.055 0.079 31.63
+FiRST (Ours) Router + LoRA 0.329 0.137 0.126 41.66
+Router 0.105 0.035 0.053 27.83
+35 Skip Decode Router + LoRA 0.028 0.011 0.016 22.30
+Router 0.024 0.010 0.014 18.32
+Random Skip Router + LoRA 0.042 0.015 0.022 32.81
+Router 0.021 0.008 0.011 23.63
+Unified Skip Router + LoRA 0.033 0.005 0.010 19.12
+Router 0.022 0.004 0.007 19.82
+FiRST (Ours) Router + LoRA 0.100 0.070 0.075 28.10
+Router 0.065 0.018 0.029 23.13
+LLaMA-3-8B
+0 Original Model Base + LoRA 0.569 0.356 0.333 82.66
+Base 0.380 0.225 0.208 68.95
+15 Skip Decode Router + LoRA 0.287 0.158 0.149 55.98
+Router 0.047 0.023 0.030 21.75
+Random Skip Router + LoRA 0.479 0.258 0.237 67.32
+Router 0.366 0.187 0.168 59.89
+Unified Skip Router + LoRA 0.466 0.250 0.230 69.58
+Router 0.273 0.134 0.122 54.57
+FiRST (Ours) Router + LoRA 0.484 0.266 0.247 68.63
+Router 0.176 0.087 0.080 42.76
+25 Skip Decode Router + LoRA 0.173 0.094 0.088 33.85
+Router 0.038 0.018 0.024 20.93
+Random Skip Router + LoRA 0.117 0.047 0.065 27.73
+Router 0.074 0.028 0.040 35.16
+Unified Skip Router + LoRA 0.349 0.158 0.142 50.59
+Router 0.177 0.074 0.068 38.74
+FiRST (Ours) Router + LoRA 0.358 0.157 0.119 56.92
+Router 0.110 0.032 0.051 25.45
+35 Skip Decode Router + LoRA 0.112 0.057 0.054 25.23
+Router 0.036 0.017 0.023 22.84
+Random Skip Router + LoRA 0.074 0.022 0.034 27.35
+Router 0.045 0.011 0.018 29.46
+Unified Skip Router + LoRA 0.075 0.021 0.034 20.25
+Router 0.039 0.011 0.017 21.24
+FiRST (Ours) Router + LoRA 0.157 0.040 0.066 26.80
+Router 0.061 0.015 0.025 22.89
+Table 7: Machine Translation Results for English to Chinese on LLaMA-3.2-3B and LLaMA-3-8B: BLEU and COMET scores for various skipping strategies.
+Layer $\downarrow$ $\alpha=0.005$ $\alpha=0.01$ $\alpha=0.025$
+R R+L R R+L R R+L
+0 0.00 0.00 0.00 0.00 0.00 0.00
+1 0.00 0.00 0.00 0.00 0.00 0.00
+2 0.00 0.00 0.00 0.00 0.00 0.00
+3 0.00 0.00 0.00 0.00 0.00 0.00
+4 0.00 0.00 0.00 0.00 0.00 0.00
+5 0.00 0.00 0.00 0.00 0.00 0.00
+6 0.00 0.00 0.00 0.00 0.00 0.00
+7 100.00 100.00 100.00 100.00 100.00 100.00
+8 100.00 100.00 100.00 100.00 100.00 100.00
+9 100.00 100.00 0.00 0.00 0.10 1.32
+10 0.00 0.00 0.00 0.00 89.25 72.67
+11 0.00 0.00 0.00 0.00 0.00 0.00
+12 0.00 0.00 100.00 100.00 100.00 100.00
+13 0.00 0.00 0.00 0.00 0.00 0.00
+14 0.00 0.00 0.00 0.59 0.00 0.25
+15 0.00 0.00 100.00 99.95 100.00 99.36
+16 0.00 0.00 0.10 2.45 100.00 99.95
+17 0.00 0.00 0.00 0.00 0.00 0.00
+18 97.79 32.24 100.00 100.00 100.00 100.00
+19 0.00 0.00 99.85 91.17 100.00 99.61
+20 0.00 0.00 100.00 100.00 100.00 99.46
+21 99.85 98.72 100.00 100.00 100.00 100.00
+22 0.00 0.00 0.00 0.00 0.00 0.00
+23 0.00 0.00 24.14 0.79 99.75 91.66
+24 0.00 0.00 0.00 0.00 0.00 0.00
+25 0.00 0.00 0.00 0.00 0.00 0.00
+26 0.00 0.00 57.31 2.45 100.00 86.02
+27 0.00 0.00 0.00 0.00 0.00 0.00
+28 0.00 0.00 0.00 0.00 0.00 0.05
+29 0.00 0.00 0.00 0.00 0.00 0.00
+30 0.00 0.00 0.00 0.00 0.00 0.00
+31 0.00 0.00 0.00 0.00 0.00 0.00
+Avg 15.55 13.47 27.54 24.92 37.16 35.95
+Table 8: Variation in skipping percentage (15–35%) with the non-skip penalization loss coefficient $\alpha$ for LLaMA-3-8B on English–German translation.
+As $\alpha$ increases, average skipping rises across both Router-only (R) and Router+LoRA (R+L) models.
 Skip (%) Model Type BERT R-1 R-L
@@
 Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
-Layer $\downarrow$ R R+L R R+L R R+L
-$\alpha\rightarrow$ 0.005 0.005 0.01 0.01 0.025 0.025
-0 0.00 0.00 0.00 0.00 0.00 0.00
-1 0.00 0.00 0.00 0.00 0.00 0.00
-2 0.00 0.00 0.00 0.00 0.00 0.00
-3 0.00 0.00 0.00 0.00 0.00 0.00
-4 0.00 0.00 0.00 0.00 0.00 0.00
-5 0.00 0.00 0.00 0.00 0.00 0.00
-6 0.00 0.00 0.00 0.00 0.00 0.00
-7 100.00 100.00 100.00 100.00 100.00 100.00
-8 100.00 100.00 100.00 100.00 100.00 100.00
-9 100.00 100.00 0.00 0.00 0.10 1.32
-10 0.00 0.00 0.00 0.00 89.25 72.67
-11 0.00 0.00 0.00 0.00 0.00 0.00
-12 0.00 0.00 100.00 100.00 100.00 100.00
-13 0.00 0.00 0.00 0.00 0.00 0.00
-14 0.00 0.00 0.00 0.59 0.00 0.25
-15 0.00 0.00 100.00 99.95 100.00 99.36
-16 0.00 0.00 0.10 2.45 100.00 99.95
-17 0.00 0.00 0.00 0.00 0.00 0.00
-18 97.79 32.24 100.00 100.00 100.00 100.00
-19 0.00 0.00 99.85 91.17 100.00 99.61
-20 0.00 0.00 100.00 100.00 100.00 99.46
-21 99.85 98.72 100.00 100.00 100.00 100.00
-22 0.00 0.00 0.00 0.00 0.00 0.00
-23 0.00 0.00 24.14 0.79 99.75 91.66
-24 0.00 0.00 0.00 0.00 0.00 0.00
-25 0.00 0.00 0.00 0.00 0.00 0.00
-26 0.00 0.00 57.31 2.45 100.00 86.02
-27 0.00 0.00 0.00 0.00 0.00 0.00
-28 0.00 0.00 0.00 0.00 0.00 0.05
-29 0.00 0.00 0.00 0.00 0.00 0.00
-30 0.00 0.00 0.00 0.00 0.00 0.00
-31 0.00 0.00 0.00 0.00 0.00 0.00
-Avg 15.55 13.47 27.54 24.92 37.16 35.95
-Table 15: Variation in skipping percentage (15-35%) with the Non-skip Penalization Loss coefficient $\alpha$ for LLaMA-3-8B on Machine Translation (English-to-German).
-As $\alpha$ increases, the skipping percentage also increases for both the Router-Only and Router-Augmented LoRA fine-tuned models.
-Similar trends are observed for other datasets as well.
+[…]
+TPOT TPOT
+Base + LoRA 0 1x 1x
+R + LoRA 15 0.90x 0.88x
+R + LoRA 25 0.82x 0.83x
+R + LoRA 35 0.69x 0.68x
+[…]
+TPOT
+Base + LoRA 0 1x
+R + LoRA 15 0.88x
+R + LoRA 20 0.81x
+R + LoRA 27 0.76x
+Table 11: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST.
+The reported values are relative to the LoRA fine-tuned base model.
+Fine-tuning improves TPOT and quality significantly.
+[…]
+TPOT TPOT
+Base + LoRA 0 1x 1x
+R + LoRA 15 0.90x 0.91x
+R + LoRA 25 0.78x 0.75x
+R + LoRA 35 0.69x 0.74x
+[…]
+TPOT
+Base + LoRA 0 1x
+R + LoRA 15 0.88x
+R + LoRA 24 0.79x
+R + LoRA 28 0.77x
+Table 12: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST.
+The reported values are relative to the LoRA fine-tuned base model.
+Fine-tuning improves TPOT and quality significantly.
+A.7 Detailed Result Table
+Tables 6 and 7 present the detailed results for the LLaMA-3.2-3B and LLaMA-3-8B models for an additional skipping percentage on Machine Translation Task.
+The results are reported using BLEU (BLEU-1, BLEU-2, BLEU) and COMET metrics, highlighting performance across different skipping percentages.
+Similarly, Table 9 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L.
+Lastly, Table 10 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
+A.8 Generalizability and Reusability of Routers
+We also investigate whether a router trained on one dataset can be applied directly to another dataset for the same task, without any additional training.
+Analysis of layer-skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere.
+To validate this idea, we train routers on English→German translation data and test them on English→Chinese translation, and likewise train on SQuAD for question answering before evaluating on the Natural Questions dataset.
+The quantitative results of these cross-dataset evaluations are presented in Table 13.
+In the QA experiments, the router trained and tested on NQ itself (denoted as Original in Table 13) and the router trained on SQuAD but tested on NQ (Cross-dataset) both skip a comparable fraction of layers.
+Despite being trained on different data, the SQuAD-trained router achieves QA accuracy (EM score) very close to that of the NQ-trained router, as detailed in Table 13.
+This indicates that the essential decision patterns learned by the router transfer well across QA datasets when using a similar level of layer skipping.
+A similar pattern emerges in machine translation.
+Whether trained on English→German or directly on English→Chinese data, routers that skip the same proportion of layers produce very similar translation quality metrics (BLEU and COMET in Table 13) on the English→Chinese task.
+This holds across different model sizes and skip settings.
+Together, these results demonstrate that routers learned on one dataset can be effectively reused on another dataset for the same task, as long as they are configured to skip a similar amount of computation.
+This reusability can lead to substantial savings in both training time and computing resources.
+Question Answering (Natural Questions)
+Setting Skip (%) EM
+LLaMA-3.2-3B
+Original 10.32 22.55
+Cross-Dataset 12.46 20.87
+Original 18.03 17.51
+Cross-Dataset 19.04 16.93
+LLaMA-3-8B
+Original 11.44 33.53
+Cross-Dataset 13.11 29.70
+Original 24.60 18.02
+Cross-Dataset 26.80 19.58
+Machine Translation (Eng–Zh)
+Setting Skip (%) BLEU COMET
+LLaMA-3.2-3B
+Original 14.69 0.37 0.55
+Cross-Dataset 16.57 0.37 0.45
+Original 26.12 0.25 0.28
+Cross-Dataset 26.16 0.25 0.30
+LLaMA-3-8B
+Original 16.46 0.25 0.43
+Cross-Dataset 14.70 0.37 0.49
+Original 27.73 0.25 0.25
+Cross-Dataset 30.00 0.25 0.24
+[…]

### v4

@@
 Let us first understand what it entails to learn an optimal subsequence of layers for any input.
-Consider the full transformer sequence to be =$\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$.
+Consider the full transformer sequence to be $\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$.
 Any optimal subsequence for an input $\mathcal{X}$: $\mathcal{F}_{OPT,\mathcal{X}}$ could be thought of as finding an optimal path through a binary tree of functions.
@@
 The autoregressive computations for the $j^{th}$ token in the input sequence (originally Eq 1), would now be modified as:
-$H^{j}_{i}$ =$=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
+$H^{j}_{i}$ $=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
 Our problem translates to navigating this binary tree to find the optimal path $\mathcal{F}_{OPT}$ for an input sequence and task.
@@
 For the first layer, it takes the tokenized input, and for each of the subsequent layers, it takes the output of the preceding layer as input.
-Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer $(i-1)$ and outputs a $B\times n\times 1$ tensor.
+Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer (i1)(i-1) and outputs a $B\times n\times 1$ tensor.
 Subsequently, corresponding to each value (or, token) in the $B\times n\times 1$ tensor, we apply a sigmoid function to ensure that all entries in the tensor are in the interval $[0,1]$.
@@
 The loss function comprises of 3 terms:
-  • Cross-entropy loss: Standard difference between actual and predicted probability distributions to ensure the quality of generation: $\mathcal{L}_{\mathrm{CE}}=-\sum_{x\in\mathcal{X}}\mathcal{Y}^{*}_{\mathcal{X}}\log(\hat{\mathcal{Y}})$.
-  • Regularization loss: Adds a penalty term to reduce overfitting to noise: $\mathcal{L}_{\mathrm{Reg}}=\sum_{i\in[m]}||R_{i}||^{2}$, where $||R_{i}||^{2}$ denotes the $\ell_{2}$ norm of the router weights for the $i^{\text{th}}$ layer router, and there are $m$ layers in the model.
+[…]
+[…]
   • Non-skip penalization loss: This is the summation of probability values across all layers of the model architecture: $\mathcal{L}_{\mathrm{PP}}=\sum_{i\in[m]}\rho_{i}$
@@
 Figure 2: Skip connection used for router training.
-With probability $p$, the sequence is processed by the layer and with probability $1-p$, the layer is skipped.
+With probability $p$, the sequence is processed by the layer and with probability 1p1-p, the layer is skipped.
 During inference, routers make the decision of whether a sequence will skip a particular layer or pass through it.
@@
 For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT.
-At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
+At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $20\%$ improvement in latency.
 Model size Model Type WMT CNN/DM SQuAD Natural Questions
@@
 It is significantly better in quality than all other baselines across all metrics for different levels of skipping.
-For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model.
+For LLaMA-3.2-3B, again FiRST is $95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model.
 Moreover, it is better than all other layer skipping strategies across all metrics.
@@
 It is significantly better in quality than all other baselines for different levels of skipping.
-For LLaMA-3.2-3B, again FiRST is $>93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model.
+For LLaMA-3.2-3B, again FiRST is $93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model.
 Moreover, it is better than all other layer skipping strategies.
@@
 WMT Summarization Question Answering
-[…]
+Split $\rightarrowDe En\rightarrowZh$ CNN/DM SQuAD NQ
 Train 3,505 8,983 3,400 3,400 3,400
@@
 A.4 Training and Inference Setup
-[…]
+  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively.
+Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization.
+The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps.
+The learning rate is set between $1e^{-4}$ and $3e^{-4}$ - a cosine scheduler is used to adjust the learning rate.
+Gradients are accumulated after five steps, and we set the regularization loss coefficient ($\lambda=0.01$), ensuring it meaningfully contributes to the overall loss without overpowering primary losses like cross-entropy and penalization loss.
+After training, we verify the router weight norms (e.g., thresholds of $0.1$ or $0.05$) to ensure they remain stable—neither exploding nor vanishing—preventing overfitting or underfitting.
+For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor ($\text{lora\_alpha}=32$).
+Our approach involves two key penalization coefficients: $\alpha$ for the router and $\beta$ for LoRA training.
+The non-skip coefficient ($\alpha$) is adjusted based on dataset characteristics, sequence length, and model depth, requiring some tuning.
+Typically, router penalization ($\alpha$) is set 2–4× higher than LoRA’s ($\beta$) based on experimental observations to maintain a consistent skipping percentage across both training phases.
+A reasonable starting point is to tune $\alpha$ for a 10–15% skipping range, then gradually increase it at regular intervals to encourage higher skipping levels.
+In addition to the above hyperparameters, we define the maximum sequence length based on the task.
+For translation, it is set to 128 for router training and 256 for LoRA training.
+Similarly, for summarization, the sequence length is set to 500 and 700, respectively.
+For Question Answering, this length is set to 512.
+The prompts for the different tasks related to training / inference are shown in Appendix A.5.
   • Joint training considerations: The decision to split training into two phases—first training the router and then the LoRA modules—was made to ensure better stability during optimization.
@@
 Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
-[…]
+Model Type $\sim$ Skipping (%) $\rightarrowDe Eng\rightarrowZh$
 TPOT TPOT
@@
 Fine-tuning improves TPOT and quality significantly.
-[…]
+Model Type $\sim$ Skipping (%) $\rightarrowDe Eng\rightarrowZh$
 TPOT TPOT
@@
 We also investigate whether a router trained on one dataset can be applied directly to another dataset for the same task, without any additional training.
-Analysis of layer-skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere.
+Analysis of layer‐skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere.
 To validate this idea, we train routers on English→German translation data and test them on English→Chinese translation, and likewise train on SQuAD for question answering before evaluating on the Natural Questions dataset.
@@
 In the QA experiments, the router trained and tested on NQ itself (denoted as Original in Table 13) and the router trained on SQuAD but tested on NQ (Cross-dataset) both skip a comparable fraction of layers.
-Despite being trained on different data, the SQuAD-trained router achieves QA accuracy (EM score) very close to that of the NQ-trained router, as detailed in Table 13.
+Despite being trained on different data, the SQuAD‐trained router achieves QA accuracy (EM score) very close to that of the NQ‐trained router, as detailed in Table 13.
 This indicates that the essential decision patterns learned by the router transfer well across QA datasets when using a similar level of layer skipping.